### v1

# A classification of Prüfer domains of integer-valued polynomials on algebras

Abstract
Let $D$ be an integrally closed domain with quotient field $K$ and $A$ a torsion-free $D$-algebra that is finitely generated as a $D$-module and such that $A\cap K=D$. We give a complete classification of those $D$ and $A$ for which the ring $\textnormal{Int}_{K}(A)=\{f\in K[X]\mid f(A)\subseteq A\}$ is a Prüfer domain. If $D$ is a semiprimitive domain, then we prove that $\textnormal{Int}_{K}(A)$ is Prüfer if and only if $A$ is commutative and isomorphic to a finite direct product of almost Dedekind domains with finite residue fields, each of them satisfying a double-boundedness condition on its ramification indices and residue field degrees.
Keywords: integer-valued polynomial, Prüfer domain, algebra, integrally closed domain, almost Dedekind domain.
MSC Primary 13F20, 13F05, 12J20, 13A18.
1 Introduction
Let $D$ be an integrally closed domain with quotient field $K$. We will always assume that $D$ is properly contained in $K$. Let $A$ be a $D$-algebra which is a torsion-free and finitely generated $D$-module. Let $B=A\otimes_{D}K$ be the extended $K$-algebra; note that $B$ is equal to the localization $(D\setminus\{0\})^{-1}A$. We can embed $K$ and $A$ inside $B$ via the natural mappings $k\mapsto 1\otimes k$ and $a\mapsto a\otimes 1$, and we will use this fact throughout this paper. For a subset $S$ of $A$ we define the ring of integer-valued polynomials over $S$ as:
$\textnormal{Int}_{K}(S,A)=\{f\in K[X]\mid f(S)\subseteq A\}.$
When $S=A$, we set $\textnormal{Int}_{K}(A,A)=\textnormal{Int}_{K}(A)$. We will also assume throughout that $A\cap K=D$ which amounts to say that $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}_{K}(D)=\textnormal{Int}(D)$, the ring of integer-valued polynomials over $D$.
Integer-valued polynomials have a long history in commutative ring theory. One problem that attracted significant interest [4, 13, 5, 17] was to classify the domains $D$ such that $\textnormal{Int}(D)$ is a Prüfer domain. Necessary and sufficient conditions on $D$ for this to occur were determined in [18]. Since the late 2000s, attention has turned to rings of the form $\textnormal{Int}_{K}(A)$ and $\textnormal{Int}_{K}(S,A)$ [33, 19, 7, 9, 10, 22, 8, 11, 14, 20, 21]. Here, as in the traditional case where $A=D$, researchers have studied situations in which these rings are or are not Prüfer. For instance, when $D=\mathbb{Z}$ and $A$ is a ring of algebraic integers, it is known that $\textnormal{Int}_{\mathbb{Q}}(A)$ is Prüfer [19, Theorem 3.7]. By contrast, if $D$ is a Dedekind domain with finite residue fields and $A$ is the matrix algebra $A=M_{n}(D)$ with $n\geq 2$, then $\textnormal{Int}_{K}(A)$ is never integrally closed [28, Corollary 3.4], hence cannot be Prüfer.
In this paper we give a complete answer to [3, Problem 28], namely, establishing exactly when $\textnormal{Int}_{K}(A)$ is a Prüfer domain. Since $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}(D)$ and overrings of Prüfer domains are Prüfer, in order for $\textnormal{Int}_{K}(A)$ to be Prüfer it is necessary that $\textnormal{Int}(D)$ be Prüfer. Recall that $D$ is an almost Dedekind domain if $D_{M}$ is a discrete valuation ring for each maximal ideal $M$ of $D$. Note that a Noetherian almost Dedekind domain is a Dedekind domain. In order for $\textnormal{Int}(D)$ to be Prüfer, it is necessary (but not sufficient) for $D$ to be an almost Dedekind domain with finite residue fields [4, Proposition 6.3]. For such a domain, we adopt the following notations and vocabulary.
Notation 1.1.
Let $D$ be an almost Dedekind domain with finite residue fields.
  • The domain $D$ contains a principal ideal domain $D_{0}$, the structure of which depends on the characteristic of $K$. If $\text{char}(K)=0$, then $D_{0}\cong\mathbb{Z}$. If $\text{char}(K)=p$, for some integral prime $p$, then $D_{0}\cong\mathbb{F}_{p}[t]$, the univariate polynomial ring over the finite field $\mathbb{F}_{p}$.
  • Let $\mathfrak{p}$ denote a nonzero prime element of $D_{0}$. Explicitly, if $D_{0}\cong\mathbb{Z}$, then $\mathfrak{p}$ is an integral prime; otherwise, $\mathfrak{p}$ is a polynomial irreducible over $\mathbb{F}_{p}$.
  • For each prime $P$ of $D$, let $v_{P}$ be the normalized $P$-adic valuation on $K$.
  • For each $\mathfrak{p}$ and each $P$ over $\mathfrak{p}$, define
    $\quad e_{P,\mathfrak{p}}=v_{P}(\mathfrak{p})\quad\text{ and }\quad f_{P,\mathfrak{p}}=[(D/P):(D_{0}/\mathfrak{p}D_{0})].$
  • For each $\mathfrak{p}$, define
    $E_{\mathfrak{p}}:=\{e_{P,\mathfrak{p}}\mid P|\mathfrak{p}D_{0}\}\quad\text{ and }\quad F_{\mathfrak{p}}:=\{f_{P,\mathfrak{p}}\mid P|\mathfrak{p}D_{0}\}.$
  • We say that $D$ satisfies the double-boundedness condition if, for each $\mathfrak{p}$, both $E_{\mathfrak{p}}$ and $F_{\mathfrak{p}}$ are bounded.
Definition 1.2.
A domain $D$ will be called an ADDB-domain (short for “almost Dedekind, doubly bounded”) if $D$ is an almost Dedekind domain with finite residue fields that satisfies the double-boundedness condition from Notation 1.1.
By the following theorem, ADDB-domains are exactly those for which $\textnormal{Int}(D)$ is Prüfer.
Theorem 1.3.
([18, Theorem 2.5]) Let $D$ be an integral domain that is not a field. Then, $\textnormal{Int}(D)$ is Prüfer if and only if $D$ is an ADDB-domain.
Thus, in order for $\textnormal{Int}_{K}(A)$ to be Prüfer, it is necessary that $D$ be an ADDB-domain. Since Prüfer domains are integrally closed, we can obtain further information on when $\textnormal{Int}_{K}(A)$ is Prüfer by studying the integral closure of $\textnormal{Int}_{K}(A)$. For this, the elements of $A$ that are integral over $D$ are prominent. The fact that $A$ is finitely generated as a $D$-module implies that $B=A\otimes_{D}K$ is finitely generated as a $K$-module. Hence, each element of $B$ is algebraic over $K$.
Definition 1.4.
We denote by $\overline{D}$ the absolute integral closure of $D$ in a fixed algebraic closure $\overline{K}$ of $K$. For an element $b\in B$, we let $\mu_{b}\in K[X]$ be the minimal monic polynomial of $b$ over $K$ and let $\Omega_{b}$ be the set of roots of $\mu_{b}$ inside $\overline{K}$. When $S\subseteq B$, we set $\Omega_{S}=\bigcup_{b\in S}\Omega_{b}$. We say that $b$ is integral over $D$ if $b$ solves a monic polynomial over $D$, we let $A^{\prime}:=\{b\in B\mid b\text{ is integral over }D\}$, and we say that $A$ is integrally closed if $A=A^{\prime}$.
Our assumptions on $A$ imply that $A\subseteq A^{\prime}$. However, there is no guarantee that $A$ is equal to $A^{\prime}$, and in fact $A^{\prime}$ may fail to be a ring because $A$ is allowed to be noncommutative (this behavior is similar to that of $D$-orders in $K$-algebras, cf. [30, Section 8]). Nevertheless, in some cases the set $A^{\prime}$ affords a description of the integral closure of $\textnormal{Int}_{K}(A)$.
Definition 1.5.
Let $\textnormal{Int}_{K}(A,A^{\prime}):=\{f\in K[X]\mid f(A)\subseteq A^{\prime}\}$. For each $n\geq 1$, let
[…]
For $\mathscr{E}\subseteq\Lambda_{n}$, let $\textnormal{Int}_{K}(\mathscr{E},\Lambda_{n}):=\{f\in K[X]\mid f(\mathscr{E})\subseteq\Lambda_{n}\}$.
Theorem 1.6.
([27, Theorem 10 & 13]) Let $D$ be an integrally closed domain. Let $n$ be the vector space dimension of $B$ over $K$. Then, $\textnormal{Int}_{K}(A,A^{\prime})$ is an integrally closed ring, and $\textnormal{Int}_{K}(A,A^{\prime})=\textnormal{Int}_{K}(\Omega_{A},\Lambda_{n})$. Moreover, if all residue rings (by nonzero ideals) of $D$ are finite, then $\textnormal{Int}_{K}(A,A^{\prime})$ is equal to the integral closure of $\textnormal{Int}_{K}(A)$.
Note that if $D$ is a Dedekind domain with finite residue fields, then Theorem 1.6 applies and $\textnormal{Int}_{K}(A,A^{\prime})$ is the integral closure of $\textnormal{Int}_{K}(A)$. It is not known whether the same is true when $D$ is a general ADDB-domain (i.e. if $D$ is not Noetherian).
We now state our main result, which says in part that $\textnormal{Int}_{K}(A)$ is Prüfer if and only if $A=A^{\prime}$. The theorem also provides structural descriptions of $A$ and $B$ that are necessary and sufficient for $\textnormal{Int}_{K}(A)$ to be Prüfer. Recall that for a $D$-module $M$, $M_{P}$ denotes the localization of $M$ at $D\setminus P$. Also, recall that a ring $R$ is semiprimitive when the Jacobson radical of $R$ is $\{0\}$.
Theorem 1.7.
Let $D$ be an ADDB-domain. Let $A$ be a $D$-algebra that is torsion-free and finitely generated as a $D$-module, and such that $A\cap K=D$. Let $B=A\otimes_{D}K$.
  1. (1)
    The following are equivalent.
    1. (i)
      $\textnormal{Int}_{K}(A)$ is Prüfer.
    2. (ii)
      $\textnormal{Int}_{K}(A)=\textnormal{Int}_{K}(A,A^{\prime})$.
    3. (iii)
      $A=A^{\prime}$.
    4. (iv)
      For all $a\in A$, $A\cap K[a]$ is integrally closed in $K[a]$.
    5. (v)
      For some $t\geq 1$, $B\cong\prod_{i=1}^{t}\mathcal{D}_{i}$, where each $\mathcal{D}_{i}$ is a division ring over $K$, and $A\cong\prod_{i=1}^{t}A_{i}$, where each $A_{i}$ is an integrally closed domain (not necessarily commutative) in $\mathcal{D}_{i}$.
    6. (vi)
      For each prime ideal $P$ of $D$, $\textnormal{Int}_{K}(A)_{P}$ is Prüfer.
    7. (vii)
      For each prime ideal $P$ of $D$, $\textnormal{Int}_{K}(A_{P})$ is Prüfer.
    8. (viii)
      For each prime ideal $P$ of $D$, $\textnormal{Int}_{K}(A_{P})$ is integrally closed.
  2. (2)
    Assume $D$ is semiprimitive. Then, $\textnormal{Int}_{K}(A)$ is Prüfer if and only if $A\cong\prod_{i=1}^{t}A_{i}$, where $A_{i}$ is the integral closure of $D$ in a finite field extension $F_{i}/K$. In this case, each $A_{i}$ is an ADDB-domain, and $A$ is a commutative ring.
Each condition in Theorem 1.7(1) implies that $\textnormal{Int}_{K}(A)$ is integrally closed. In analogy with condition (viii), we suspect that $\textnormal{Int}_{K}(A)$ is Prüfer if and only if $\textnormal{Int}_{K}(A)$ is integrally closed; however, we have only been able to prove this with additional assumptions on $D$ or $\textnormal{Int}_{K}(A)$.
Conjecture 1.8.
Let $D$ and $A$ be as in the statement of Theorem 1.7. If $\textnormal{Int}_{K}(A)$ is integrally closed, then $\textnormal{Int}_{K}(A)$ is Prüfer.
The full resolution of Conjecture 1.8 remains an open problem. We can show that the conjecture is true if $D$ is an ADDB-domain that is also a Dedekind domain or, more generally, if $\textnormal{Int}_{K}(A)$ is well-behaved with respect to localization at primes of $D$:
Theorem 1.9.
Let $D$ and $A$ be as in the statement of Theorem 1.7. If $\textnormal{Int}_{K}(A)_{P}=\textnormal{Int}_{K}(A_{P})$ for all primes $P\subset D$, then Conjecture 1.8 is true.
In particular, Theorem 1.9 holds when $D$ is Dedekind, because in that case $\textnormal{Int}_{K}(A)_{P}=\textnormal{Int}_{K}(A_{P})$ by [29, Lemma 3.2].
The paper is organized as follows. In Section 2, we classify the finite subsets $S\subseteq A$ such that $\textnormal{Int}_{K}(S,A)$ is Prüfer. This gives us some control on the overrings of $\textnormal{Int}_{K}(A)$, and these are used in Section 3 where we derive a number of necessary conditions on $A$ in order for $\textnormal{Int}_{K}(A)$ to be Prüfer. In Theorem 3.4, we prove that if $\textnormal{Int}_{K}(A)$ is Prüfer, then $A$ and $B$ have the structure specified in condition (v) of Theorem 1.7(1). Moreover, we prove Theorem 1.7(2) (see Corollary 3.13), which shows that if $D$ is semiprimitive, then $A$ must be commutative. If $D$ is not semiprimitive, then it is possible that $\textnormal{Int}_{K}(A)$ is Prüfer and $A$ is noncommutative; an example of such an algebra $A$ is given in Section 5 at the end of the paper. We complete the proof of Theorem 1.7 in Section 4 by constructing a family of Prüfer domains contained in $K[X]$. We show that $\textnormal{Int}_{K}(A,A^{\prime})$ must contain one of these Prüfer domains, and hence is itself Prüfer (Corollary 4.12). Under any of conditions (iii) – (viii) in Theorem 1.7(1), we demonstrate that $\textnormal{Int}_{K}(A)=\textnormal{Int}_{K}(A,A^{\prime})$, and conclude $\textnormal{Int}_{K}(A)$ is Prüfer. Section 4 closes with a discussion of Conjecture 1.8 and Theorem 1.9.
2 Finite subsets of AA
Our first goal is to determine when $\textnormal{Int}_{K}(S,A)$ is Prüfer for a finite subset $S\subseteq A$. Since Prüfer domains are integrally closed, we begin by studying the integral closure of $\textnormal{Int}_{K}(S,A)$. In [23], subsets of $\overline{D}$ were used to describe the integral closure of $\textnormal{Int}_{K}(S,A)$. Recall from Definition 1.4 that for $b\in B$ and $S\subseteq B$, we let $\Omega_{b}\subseteq\overline{K}$ be the set of roots of $\mu_{b}$ in $\overline{K}$, and $\Omega_{S}=\bigcup_{b\in S}\Omega_{b}$.
Theorem 2.1.
  1. (1)
    ([23, Proposition 4.2]) Let $\Omega$ be a finite subset of $\overline{D}$. Then, $\textnormal{Int}_{K}(\Omega,\overline{D})$ is Prüfer if and only if $D$ is Prüfer.
  2. (2)
    ([23, Corollary 1.5]) Let $S$ be a finite subset of $A$. Then, the integral closure of $\textnormal{Int}_{K}(S,A)$ is equal to $\textnormal{Int}_{K}(\Omega_{S},\overline{D})=\{f\in K[X]\mid f(\Omega_{S})\subseteq\overline{D}\}$. Thus, $\textnormal{Int}_{K}(S,A)$ is Prüfer if and only if $D$ is Prüfer and $\textnormal{Int}_{K}(S,A)$ is integrally closed.
It remains to determine conditions on $S$ and $A$ under which $\textnormal{Int}_{K}(S,A)$ is integrally closed. We begin by considering the case where $S=\{a\}$ for some $a\in A$. Even when $S$ is a singleton set, it is not always apparent whether or not $\textnormal{Int}_{K}(S,A)$ will be integrally closed. Consider the following two examples involving $2\times 2$ matrices over $\mathbb{Z}$.
Example 2.2.
Let $D=\mathbb{Z}$, $A=M_{2}(\mathbb{Z})$, and $\mu(X)=X^{2}-2X-4$, with roots $1\pm\sqrt{5}$. Let $\theta=\tfrac{1+\sqrt{5}}{2}$. Let $a=[\begin{smallmatrix}0&4\\ 1&2\end{smallmatrix}]$ and $b=[\begin{smallmatrix}0&2\\ 2&2\end{smallmatrix}]$, both of which are matrices in $A$ with minimal polynomial $\mu$. Since $a$ and $b$ share the same minimal polynomial, they are conjugate over $GL(2,\mathbb{Q})$, although one may check that they are not conjugate over $GL(2,\mathbb{Z})$. We will show that $\textnormal{Int}_{\mathbb{Q}}(\{a\},A)$ is not integrally closed, but $\textnormal{Int}_{\mathbb{Q}}(\{b\},A)$ is integrally closed.
Let $\mathcal{R}=\textnormal{Int}_{\mathbb{Q}}(\{1\pm\sqrt{5}\},\mathbb{Z}[\theta])$. Then, $\mathcal{R}$ is the integral closure of both $\textnormal{Int}_{\mathbb{Q}}(\{a\},A)$ and $\textnormal{Int}_{\mathbb{Q}}(\{b\},A)$ by Theorem 2.1. To see that $\textnormal{Int}_{\mathbb{Q}}(\{a\},A)$ is not integrally closed, note that $\tfrac{X}{2}\in\mathcal{R}$, but $\tfrac{a}{2}\notin A$. Hence, $\textnormal{Int}_{\mathbb{Q}}(\{a\},A)\subsetneqq\mathcal{R}$.
Next, we show that $\textnormal{Int}_{\mathbb{Q}}(\{b\},A)=\mathcal{R}$. We have $\textnormal{Int}_{\mathbb{Q}}(\{b\},A)\subseteq\mathcal{R}$ by Theorem 2.1. For the reverse containment, let $f\in\mathcal{R}$. Then, $f(1+\sqrt{5})\in\mathbb{Z}[\theta]$, so there exist $c_{0},c_{1}\in\mathbb{Z}$ such that $f(1+\sqrt{5})=c_{0}+c_{1}\theta$. Note that $\mathbb{Q}(b)\cong\mathbb{Q}[X]/(\mu(X))\cong\mathbb{Q}(1+\sqrt{5})$, so there exists a ring isomorphism $\phi:\mathbb{Q}(1+\sqrt{5})\to\mathbb{Q}(b)$ such that $\phi(1+\sqrt{5})=b$. As $\phi(\theta)=\tfrac{b}{2}=[\begin{smallmatrix}0&1\\ 1&1\end{smallmatrix}]$ and polynomials in $\mathbb{Q}[X]$ commute with $\phi$, we have
$f(b)=f(\phi(1+\sqrt{5}))=\phi(f(1+\sqrt{5}))=c_{0}+c_{1}[\begin{smallmatrix}0&1\\ 1&1\end{smallmatrix}]\in A.$
This shows that $f\in\textnormal{Int}_{\mathbb{Q}}(\{b\},A)$. Thus, $\textnormal{Int}_{\mathbb{Q}}(\{b\},A)=\mathcal{R}$ is integrally closed.
Example 2.3.
Again let $D=\mathbb{Z}$ and $A=M_{2}(\mathbb{Z})$. Let $k\in\mathbb{Z}$, $k>0$. Let $a=[\begin{smallmatrix}0&k\\ k&0\end{smallmatrix}]$ and $b=[\begin{smallmatrix}k&0\\ 0&-k\end{smallmatrix}]$. Then, $\mu(X)=X^{2}-k^{2}$ is the minimal polynomial of both $a$ and $b$. As in the previous example, $a$ and $b$ are conjugate over $GL(2,\mathbb{Q})$, but are not conjugate over $GL(2,\mathbb{Z})$. We demonstrate that $\textnormal{Int}_{\mathbb{Q}}(\{a\},A)$ is not integrally closed, but $\textnormal{Int}_{\mathbb{Q}}(\{b\},A)$ is integrally closed.
This time, the integral closure of both $\textnormal{Int}_{\mathbb{Q}}(\{a\},A)$ and $\textnormal{Int}_{\mathbb{Q}}(\{b\},A)$ is $\mathcal{R}=\textnormal{Int}_{\mathbb{Q}}(\{\pm k\},\mathbb{Z})$. We have $\tfrac{X-k}{2k}\in\mathcal{R}\setminus\textnormal{Int}_{\mathbb{Q}}(\{a\},A)$, because
$\tfrac{1}{2k}(a-k)=\begin{bmatrix}-k/2&k/2\\ k/2&-k/2\end{bmatrix}\notin A.$
Thus, $\textnormal{Int}_{\mathbb{Q}}(\{a\},A)$ is not integrally closed. However, since $b$ is a diagonal matrix, for each $f\in\mathcal{R}$ we have $f(b)=[\begin{smallmatrix}f(k)&0\\ 0&f(-k)\end{smallmatrix}]$, and this is in $A$ because $f(\pm k)\in\mathbb{Z}$. Hence, $\textnormal{Int}_{\mathbb{Q}}(\{b\},A)=\mathcal{R}$.
One distinguishing feature between $a$ and $b$ in these examples is the structure of the algebras $A\cap K[a]$ and $A\cap K[b]$. For each $a\in A$, $K[a]$ is a commutative $K$-subalgebra of $B$ and the intersection $A\cap K[a]$ is a $D$-subalgebra of $K[a]$. Recall that a commutative ring $R$ (possibly with zero divisors) is said to be integrally closed if it is integrally closed in its total ring of fractions $T$ (i.e. every element of $T$ which satisfies a monic polynomial over $R$ belongs to $R$).
Lemma 2.4.
Let $a\in A$, let $R=A\cap K[a]$, and let $T$ be the total ring of fractions of $R$.
  1. (1)
    $T=K[a]$.
  2. (2)
    $R$ is integrally closed if and only if $R$ is equal to the integral closure of $D$ in $K[a]$.
Proof.
(1) Recall that $\mu_{a}\in D[X]$ is the minimal polynomial of $a$ over $D$. Factor $\mu_{a}$ over $K$ as $\mu_{a}(X)=\prod_{i=1}^{t}p_{i}(X)^{e_{i}}$, where each $p_{i}\in K[x]$ is irreducible and each $e_{i}\geq 1$. Then,
$K[a]\cong K[x]/(\mu_{a}(X))\cong\prod_{i=1}^{t}K[x]/(p_{i}(X)^{e_{i}}).$
Each ring $K[x]/(p_{i}(X)^{e_{i}})$ is local with maximal ideal generated by $p_{i}(X)$. In each such ring, the set of zero divisors is equal to the maximal ideal, and all other elements are units. Hence, each of these rings is equal to its own total ring of fractions. It follows that $K[a]$ is also equal to its own total ring of fractions. Now, $D[a]\subseteq R\subseteq K[a]$, and considering the total rings of fractions of these rings, we see that $K[a]\subseteq T\subseteq K[a]$. Thus, $T=K[a]$.
(2) Let $R^{\prime}$ be the integral closure of $R$ in $T$, and let $D^{\prime}$ be the integral closure of $D$ in $K[a]$. We have $D\subseteq R$, and $T=K[a]$ by (1), so $D^{\prime}\subseteq R^{\prime}$. Conversely, let $b\in R^{\prime}\subseteq K[a]$. Then, there exists a monic $f\in R[X]\subseteq A[X]$ such that $f(b)=0$. Since each coefficient of $f$ is in $A$ and $A$ is integral over $D$, $b$ is integral over $D$. Hence, $b\in D^{\prime}$ and $R^{\prime}\subseteq D^{\prime}$. ∎
In both Example 2.2 and Example 2.3, the ring $A\cap K[a]$ is not integrally closed. In the former example, $a/2\in K[A]\setminus A$ is integral over $D$, and in the latter $(a-k)/2k\in K[a]\setminus A$ is integral over $D$. As we now proceed to show, $\textnormal{Int}_{K}(\{a\},A)$ is integrally closed if and only if $A\cap K[a]$ is integrally closed.
Lemma 2.5.
Let $b\in B$.
  1. (1)
    $b$ is integral over $D$ if and only if $\Omega_{b}\subseteq\overline{D}$.
  2. (2)
    For each $f\in K[X]$, we have $\Omega_{f(b)}=f(\Omega_{b})$.
Proof.
Let $n$ be the vector space dimension of $B$ over $K$. Then, $B$ embeds into the matrix ring $M_{n}(K)$, and under this embedding $K$ corresponds to the field of scalar matrices in $M_{n}(K)$. Thus, each $b\in B$ corresponds to a matrix $M_{b}\in M_{n}(K)$ with minimal polynomial $\mu_{b}\in K[X]$. The eigenvalues of $M_{b}$ are exactly the roots of $\mu_{b}$, and it is shown in [27, Lemma 2] that $\mu_{b}\in D[X]$ if and only if each eigenvalue of $M_{b}$ is in $\overline{D}$. This proves (1). For (2), we use the fact that for a matrix $C\in M_{n}(K)$, a scalar $\lambda\in K$, and a polynomial $f\in K[X]$, $\lambda$ is an eigenvalue of $C$ if and only if $f(\lambda)$ is an eigenvalue of $f(C)$. ∎
Proposition 2.6.
Let $a\in A$. The following are equivalent.
  1. (1)
    $\textnormal{Int}_{K}(\{a\},A)$ is integrally closed.
  2. (2)
    $\textnormal{Int}_{K}(\{a\},A)=\textnormal{Int}_{K}(\Omega_{a},\overline{D})$.
  3. (3)
    $A\cap K[a]$ is integrally closed.
Proof.
(1) $\Leftrightarrow$ (2) This follows from Theorem 2.1, because $\{a\}$ is finite.
(3) $\Rightarrow$ (2) Assume that $A\cap K[a]$ is integrally closed. From Theorem 2.1, $\textnormal{Int}_{K}(\Omega_{a},\overline{D})$ is the integral closure of $\textnormal{Int}_{K}(\{a\},A)$, and so it is always true that $\textnormal{Int}_{K}(\{a\},A)\subseteq\textnormal{Int}_{K}(\Omega_{a},\overline{D})$. Let $f\in\textnormal{Int}_{K}(\Omega_{a},\overline{D})$. Then, $f(a)\in K[a]$, because $f\in K[X]$, and $f(a)$ is integral over $D$ by Lemma 2.5 because $\Omega_{f(a)}=f(\Omega_{a})\subseteq\overline{D}$ by Lemma 2.5. By Lemma 2.4, $A\cap K[a]$ is the integral closure of $D$ in $K[a]$. Thus, $f(a)\in A\cap K[a]\subseteq A$, $f\in\textnormal{Int}_{K}(\{a\},A)$, and $\textnormal{Int}_{K}(\{a\},A)=\textnormal{Int}_{K}(\Omega_{a},\overline{D})$.
(1) $\Rightarrow$ (3) Assume that $\textnormal{Int}_{K}(\{a\},A)$ is integrally closed. Let $b\in K[a]$ be integral over $D$. Then, $b=f(a)$ for some $f\in K[X]$, and there exists a monic $g\in D[X]$ such that $g(b)=0$. So, $g\circ f\in\textnormal{Int}_{K}(\{a\},A)$. Since $g\in D[X]\subseteq\textnormal{Int}_{K}(\{a\},A)$, this implies that $f$ is integral over $\textnormal{Int}_{K}(\{a\},A)$, and hence $f\in\textnormal{Int}_{K}(\{a\},A)$. Thus, $b=f(a)\in A\cap K[a]$. This shows that $A\cap K[a]$ is the integral closure of $D$ in $K[a]$, so $A\cap K[a]$ is integrally closed by Lemma 2.4. ∎
We now have necessary and sufficient conditions on $A$ and $D$ to describe when $\textnormal{Int}_{K}(\{a\},A)$ is Prüfer. The conditions extend naturally to the case where $S\subseteq A$ is finite.
Theorem 2.7.
Let $a\in A$. The following are equivalent.
  1. (1)
    $\textnormal{Int}_{K}(\{a\},A)$ is Prüfer.
  2. (2)
    $D$ is Prüfer and $\textnormal{Int}_{K}(\{a\},A)$ is integrally closed.
  3. (3)
    $D$ is Prüfer and $A\cap K[a]$ is integrally closed.
Proof.
Apply Theorem 2.1 and Proposition 2.6. ∎
Corollary 2.8.
Let $D$ be a Prüfer domain and let $S\subset A$ be a finite subset. Then $\textnormal{Int}_{K}(S,A)$ is Prüfer if and only if for each $a\in S$, $\textnormal{Int}_{K}(\{a\},A)$ is Prüfer.
Proof.
We have just to observe that
$\textnormal{Int}_{K}(S,A)=\bigcap_{a\in S}\textnormal{Int}_{K}(\{a\},A).$
If $\textnormal{Int}_{K}(S,A)$ is Prüfer then each of its overrings $\textnormal{Int}_{K}(\{a\},A)$ is Prüfer as well. Conversely, if $\textnormal{Int}_{K}(\{a\},A)$ is Prüfer, then by Theorem 2.1, $\textnormal{Int}_{K}(\{a\},A)=\textnormal{Int}_{K}(\Omega_{a},\overline{D})$. The intersection of these rings over all $a\in S$ is equal to $\textnormal{Int}_{K}(\Omega_{S},\overline{D})$ which is Prüfer because $\Omega_{S}=\bigcup_{a\in S}\Omega_{a}$ is finite. ∎
3 Necessary conditions
In this section, we will establish necessary conditions for $\textnormal{Int}_{K}(A)$ to be Prüfer. Since one of our standard assumptions on $A$ is that $A\cap K=D$, we have $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}(D)$. Thus, if $\textnormal{Int}_{K}(A)$ is Prüfer, then so is $\textnormal{Int}(D)$. If the reader wishes, for the remainder of this section they may consider $D$ to be an ADDB-domain (i.e. $D$ is such that $\textnormal{Int}(D)$ is Prüfer), since this is the assumption we will make in our major theorems. However, we state our intermediary results using only the properties of the integrally closed domain $D$ that are needed to reach the desired conclusion.
Recall that an element $b\in B$ is integral over $D$ if $b$ solves a monic polynomial in $D[X]$, and that $A$ is integrally closed if $A=A^{\prime}=\{b\in B\mid b\text{ is integral over }D\}$. While $A^{\prime}$ may not be a ring, $A\cap K[a]$ is a commutative subring of $K[a]$ for each $a\in A$. So, the integral closure of $A\cap K[a]$ is a commutative ring, and these rings allow us to describe when $A=A^{\prime}$.
Lemma 3.1.
The ring $A$ is integrally closed if and only if, for each $a\in A$, $A\cap K[a]$ is integrally closed.
Proof.
Clearly, if $A=A^{\prime}$ and $a\in A$, then every element $a^{\prime}$ of $K[a]$ that is integral over $A\cap K[a]$ is also integral over $D$. Hence, $a^{\prime}\in A$ and $A\cap K[a]$ is integrally closed. Conversely, suppose that $A\cap K[a]$ is integrally closed for each $a\in A$ and let $a^{\prime}\in A^{\prime}$. Since $a^{\prime}\in B=(D\setminus\{0\})^{-1}A$, there exist $a\in A$ and $d\in D$ such that $a^{\prime}=\frac{a}{d}$. So, $a^{\prime}\in K[a]$ and is integral over $D\subset D[a]\subseteq A\cap K[a]$. It follows that $a^{\prime}\in A\cap K[a]\subset A$. ∎
Recall that a ring $R$ is reduced if it has no nonzero nilpotent elements. One obstruction to $\textnormal{Int}_{K}(A)$ being Prüfer is the presence of nonzero nilpotent elements of $A$. In the next two lemmas, we show that if $\textnormal{Int}_{K}(A)$ is Prüfer, then $A$ must be reduced. We thank Hwankoo Kim for pointing out Lemma 3.2.
Lemma 3.2.
The $D$-topology on $A$ is Hausdorff. That is, $\bigcap_{d\in D\setminus\{0\}}dA=\{0\}$.
Proof.
Let $I=\bigcap_{d\in D\setminus\{0\}}dA$. Then, $I$ is a divisible $D$-submodule of $A$ because $dI=I$ for all nonzero $d\in D$. Hence, $I$ is injective because $A$ is torsion-free [15, Theorem 2.4.5(3)], and so $I$ is a direct summand of $A$. Thus, $I$ is a finitely generated $D$-submodule of $A$. Let $d\in D$ such that $d$ is not a unit of $D$, which exists because $D$ is not a field. By [1, Corollary 2.5], there exists $c\in D$ such that $(1+cd)I=\{0\}$. Since $d$ is not a unit of $D$, $1+cd\neq 0$. We conclude that $I=\{0\}$. ∎
Lemma 3.3.
  1. (1)
    Let $a\in A$. If $A\cap K[a]$ is integrally closed, then $K[a]$ is reduced.
  2. (2)
    If $\textnormal{Int}_{K}(A)$ is Prüfer, then both $A$ and $B$ are reduced.
Proof.
(1) We prove the contrapositive. Suppose that $K[a]$ is not reduced. Then, there exists a nonzero $b\in K[a]$ and $n\geq 2$ such that $b^{n}=0$. Let $f\in K[X]$ be such that $b=f(a)$. Write $f(X)=g(X)/d$, where $g\in D[X]$ and $d\in D$. Since $f(a)\neq 0$, $g(a)$ is also nonzero. So, there exists a nonzero $c\in D$ such that $g(a)\notin cA$ by Lemma 3.2. Then, the polynomial $g(x)/c\notin\textnormal{Int}_{K}(\{a\},A)$. However, $f(a)^{n}=0$, so
$\Big{(}\dfrac{g(X)}{c}\Big{)}^{n}=\Big{(}\dfrac{df(X)}{c}\Big{)}^{n}\in\textnormal{Int}_{K}(\{a\},A).$
Thus, $\textnormal{Int}_{K}(\{a\},A)$ is not integrally closed. By Proposition 2.6, $A\cap K[a]$ is not integrally closed.
(2) Suppose that $\textnormal{Int}_{K}(A)$ is Prüfer, but $A$ is not reduced. Let $a\in A$ be nonzero and nilpotent. Since $\textnormal{Int}_{K}(A)$ is Prüfer, its overring $\textnormal{Int}_{K}(\{a\},A)$ is also Prüfer, and hence is integrally closed. But, $K[a]$ is not reduced, so by (1), $A\cap K[a]$ is not integrally closed. By Proposition 2.6, $\textnormal{Int}_{K}(\{a\},A)$ is not integrally closed, which is a contradiction. Thus, $A$ must be reduced, and it follows that $B$ is also reduced. ∎
Theorem 3.4.
Suppose that $\textnormal{Int}_{K}(A)$ is a Prüfer domain. Then, $A=A^{\prime}$, and $B$ and $A$ have the following structure:
  • For some $t\geq 1$, $B\cong\prod_{i=1}^{t}\mathcal{D}_{i}$, where $\mathcal{D}_{i}$ is a division ring over $K$ for each $i=1,\ldots,t$, and
  • $A\cong\prod_{i=1}^{t}A_{i}$, where $A_{i}\subset\mathcal{D}_{i}$ is an integrally closed domain (but is not necessarily commutative) for each $i=1,\ldots,t$.
Proof.
Since $\textnormal{Int}_{K}(A)$ is Prüfer, each overring of the form $\textnormal{Int}_{K}(\{a\},A)$, where $a\in A$, is also Prüfer. By Theorem 2.7, $A\cap K[a]$ is integrally closed for each $a\in A$. So, $A=A^{\prime}$ by Lemma 3.1. Also, both $A$ and $B$ are reduced by Lemma 3.3.
Now, since $B$ has finite dimension over $K$ it is artinian (see e.g. [6, §23, p. 158]), and $B$ contains no nonzero nilpotent ideal because it is reduced. By Wedderburn’s theorems [30, Theorem 7.1, Theorem 7.4], $B$ is semisimple, thus isomorphic to a finite direct product of matrix rings $M_{n_{i}}(\mathcal{D}_{i})$, where $\mathcal{D}_{i}$ is a division ring over (an isomorphic copy of) $K$ for each $i$. If any $n_{i}\geq 2$, then $M_{n_{i}}(\mathcal{D}_{i})$—and hence BB—contains nonzero nilpotent elements, which is impossible because $B$ is reduced. Hence, $B\cong\prod_{i=1}^{t}\mathcal{D}_{i}$ for some $t\geq 1$.
For the structure of $A$, let $e_{i}=1_{\mathcal{D}_{i}}$ for each $1\leq i\leq t$. Then, $\{e_{1},\ldots,e_{t}\}$ is the set of orthogonal idempotents related to the above decomposition of $B$. Each $e_{i}$ is a root of $X^{2}-X\in D[X]$, and so is in $A$ because $A=A^{\prime}$. Thus, $A$ decomposes as $\prod_{i=1}^{t}A_{i}$, where $A_{i}=Ae_{i}\subset Be_{i}=\mathcal{D}_{i}$; in particular, $A_{i}$ is a (possibly noncommutative) domain for each $i$. Finally, assume $b_{i}^{\prime}\in\mathcal{D}_{i}$ is integral over $De_{i}\cong D$ with minimal polynomial $\mu_{i}(X)\in D[X]$. Then, the element $b=(b_{1},\ldots,b_{t})\in B$ with $b_{i}=b_{i}^{\prime}$ and $b_{j}=0$ for $j\neq i$ is a root of $X\mu_{i}(X)$, and hence is integral over $D$. Since $A=A^{\prime}$, it follows that $b_{i}^{\prime}\in A_{i}$, and so $A_{i}=A_{i}^{\prime}$. ∎
For any ring $R$, let $\mathscr{J}(R)$ be the Jacobson radical of $R$. We recall that $R$ is semiprimitive if $\mathscr{J}(R)=\{0\}$. If $D$ is semiprimitive (which occurs, for instance, when $D=\mathbb{Z}$), we can impose even stronger necessary conditions on $A$ in order for $\textnormal{Int}_{K}(A)$ to be Prüfer. To do this, we first show that in some cases the existence of a nonzero nilpotent element in a simple residue ring of $A$ is enough to force $\textnormal{Int}_{K}(A)$ to not be Prüfer.
Lemma 3.5.
Assume that $D$ and $A$ satisfy both conditions (i) and (ii) below:
  1. (i)
    $D$ contains a principal ideal $\pi D$ of finite index.
  2. (ii)
    $A/\pi A$ has a simple residue ring isomorphic to $M_{n}(F)$, where $n\geq 2$ and $F$ is a finite field.
Then, the following hold.
  1. (1)
    There exists $a\in A$ such that $a^{2}\in\pi^{2}A$, but $a\notin\pi A$.
  2. (2)
    $\textnormal{Int}_{K}(A)$ is not Prüfer.
Proof.
(1) Let $R_{1}=A/\pi A$ and $R_{2}=A/\pi^{2}A$, which have Jacobson radicals $J_{1}:=\mathscr{J}(R_{1})$ and $J_{2}:=\mathscr{J}(R_{2})$. By (i), $D/\pi D$ is finite, and $A$ is finitely generated as a $D$-module, so $R_{1}$ is finite. So, $R_{2}/J_{2}\cong R_{1}/J_{1}$ is a finite semisimple ring. By (ii), $R_{2}/J_{2}$ contains a direct summand $S\cong M_{n}(F)$.
Write $1_{S}=\overline{e_{1}}+\cdots+\overline{e_{n}}$ for primitive orthogonal idempotents $\overline{e_{1}},\ldots,\overline{e_{n}}\in S$. Then, $\overline{e_{1}}S\overline{e_{2}}\neq 0$ in $R_{2}/J_{2}$, and we may lift $\overline{e_{1}}$ and $\overline{e_{2}}$ to $R_{2}$ (see e.g. [16, Prop. 21.25]). Thus, $R_{2}$ contains orthogonal idempotents $e_{1}$ and $e_{2}$ such that $e_{1}R_{2}e_{2}\not\subseteq J_{2}$. Hence, there exists $b\in e_{1}R_{2}e_{2}$ such that $b\notin\pi R_{2}\subseteq J_{2}$. However, $b^{2}=0$ because $(e_{1}R_{2}e_{2})(e_{1}R_{2}e_{2})=\{0\}$. It follows that the desired element $a$ exists in $A$.
(2) Let $a\in A$ satisfy the conclusion of (1). Then, $\textnormal{Int}_{K}(\{a\},A)$ is not integrally closed, because $X^{2}/\pi^{2}\in\textnormal{Int}_{K}(\{a\},A)$, but $X/\pi\notin\textnormal{Int}_{K}(\{a\},A)$. Since $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}_{K}(\{a\},A)$, this implies that $\textnormal{Int}_{K}(A)$ is not Prüfer. ∎
Theorem 3.6.
Let $D$ be a Prüfer domain and assume that $A$ contains a maximal ideal $\mathfrak{m}$ such that $A/\mathfrak{m}\cong M_{n}(F)$, where $n\geq 2$ and $F$ is a finite field. Then, $\textnormal{Int}_{K}(A)$ is not Prüfer.
Proof.
Let $P=D\cap\mathfrak{m}$. We first verify that $P$ is a prime ideal of $D$. Certainly, $P$ is a proper ideal of $D$ because $1\notin P$. Let $cd\in P$. Then, $cd\in\mathfrak{m}$, and since both $c$ and $d$ are central in $A$, we have $cAd\subseteq\mathfrak{m}$. The ring $A$ is noncommutative, but maximal ideals in noncommutative rings are still prime [16, p. 165], so the condition $cAd\subseteq\mathfrak{m}$ implies that either $c\in\mathfrak{m}$ or $d\in\mathfrak{m}$ [16, Prop. 10.2]. Thus, either $c$ or $d$ is in $P$.
Next, recall that $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}(D)$. If $P$ has infinite index in $D$, then by [2, Proposition I.3.4] we have $\textnormal{Int}(D)\subseteq D_{P}[X]$ and so $\textnormal{Int}_{K}(A)$ cannot be Prüfer. So, assume $D:P]<\infty$ and consider the $D_{P}$-algebra $A_{P}=(D\setminus P)^{-1}A$. Since $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}_{K}(A_{P})$, it suffices to show that $\textnormal{Int}_{K}(A_{P})$ is not Prüfer. Because $D$ is Prüfer, $D_{P}$ is a valuation domain. If $PD_{P}$ is not principal, then by [2, Proposition I.3.16] we have that $\textnormal{Int}(D_{P})=D_{P}[X]$, so the subring $\textnormal{Int}_{K}(A_{P})$ cannot be Prüfer. Suppose now that $PD_{P}$ is principal and let $\pi\in D_{P}$ be a generator of $PD_{P}$ Then, $A_{P}/\pi A_{P}$ has $M_{n}(F)$ as a residue ring and $D_{P}:\pi D_{P}]=[D:P]$ is finite. By Lemma 3.5, $\textnormal{Int}_{K}(A_{P})$ is not Prüfer, and we are done. ∎
Corollary 3.7.
If $\textnormal{Int}_{K}(A)$ is Prüfer, then every simple residue ring of $A$ is a finite field.
Proof.
Assume that $\textnormal{Int}_{K}(A)$ is Prüfer. Then, $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}(D)$ because $A\cap K=D$, and so $\textnormal{Int}(D)$ is Prüfer. By [4, Prop. 6.3], $D$ must be an almost Dedekind domain with finite residue fields, hence $D$ is a Prüfer domain. Since $\textnormal{Int}_{K}(A)$ is Prüfer, the condition of Theorem 3.6 cannot hold. Therefore, every simple residue ring of AA must be a finite field. ∎
When $D$ is semiprimitive, combining Theorem 3.4 with Corollary 3.7 allows us to show that $A$ is a commutative ring whenever $\textnormal{Int}_{K}(A)$ is Prüfer. In that case, the structure of $A$ is further restricted by Theorem 1.3. Towards this end, we now study how $D$ being semiprimitive affects $A$.
In what follows, for any ring $R$, let $\textnormal{Max}(R)$ be the set of maximal ideals of $R$, called the maximal spectrum of $R$.
Lemma 3.8.
Assume that $D$ is one-dimensional.
  1. (1)
    Let $\mathfrak{m}\in\textnormal{Max}(A)$. Then, $\mathfrak{m}\cap D\in\textnormal{Max}(D)$.
  2. (2)
    Let $M\in\textnormal{Max}(D)$. Then, there exists a maximal ideal of $A$ containing $MA$.
Proof.
(1) Let $M=\mathfrak{m}\cap D$. Working as in the proof of Theorem 3.6, we see that $M$ is a prime ideal of $D$. It remains to show that $M\neq\{0\}$. If this were the case, then $D$ would be isomorphic to a subring of $A/\mathfrak{m}$, because
$D=D/(\mathfrak{m}\cap D)\cong(D+\mathfrak{m})/\mathfrak{m}\subseteq A/\mathfrak{m}.$
However, $A/\mathfrak{m}$ is a simple ring and $D$ is central in $A$, so this would force $D$ to be a field, which is a contradiction. Thus, $M$ is a nonzero prime ideal of the one-dimensional domain $D$, and hence $M$ is maximal.
(2) Consider the localization $D_{M}$ and the $D_{M}$-algebra $A_{M}$. The ring $A_{M}$ cannot be simple, since then $dA_{M}=A_{M}$ for each $d\in D\setminus\{0\}$. This implies that $K=A_{M}\cap K$, and consequently $D\subsetneqq A\cap K$. So, $A_{M}$ contains a nonzero maximal ideal $\mathfrak{M}$. The ideal $\mathfrak{M}\cap D_{M}$ is maximal in $D_{M}$, so we have $MD_{M}=\mathfrak{M}\cap D_{M}$. It follows that the contraction of $\mathfrak{M}$ to $A$ is a proper ideal of $A$ that contains $M$. ∎
Proposition 3.9.
Assume that $B$ is a division ring.
  1. (1)
    If $\mathfrak{I}$ is a two-sided ideal of $A$ such that $\mathfrak{I}\cap D=\{0\}$, then $\mathfrak{I}=\{0\}$.
  2. (2)
    If $D$ is one-dimensional and semiprimitive, then the intersection of all the maximal ideals of $A$ is $\{0\}$. Consequently, $A$ is semiprimitive.
Proof.
(1) Assume that $\mathfrak{I}$ is a two-sided ideal of $A$ with $\mathfrak{I}\cap D=\{0\}$. Let $U=D\setminus\{0\}$. Since $B=U^{-1}A$ is a division ring, the extended ideal $U^{-1}\mathfrak{I}$ of $B$ is equal to either $\{0\}$ or $B$. If $U^{-1}\mathfrak{I}=B$, then $1=d^{-1}d$ for some nonzero $d\in\mathfrak{I}\cap D$, which is a contradiction. Thus, $U^{-1}\mathfrak{I}=\{0\}$, and this implies that $\mathfrak{I}=\{0\}$.
(2) Assume that $D$ is semiprimitive and let $\mathfrak{I}=\bigcap_{\mathfrak{m}\in\textnormal{Max}(A)}\mathfrak{m}$. Let $d\in\mathfrak{I}\cap D$. Using Lemma 3.8 and the fact that $D$ is commutative and semiprimitive,
$d\in\bigcap_{\mathfrak{m}\in\textnormal{Max}(A)}(\mathfrak{m}\cap D)=\bigcap_{M\in\textnormal{Max}(D)}M=\{0\}.$
Thus, $\mathfrak{I}\cap D=\{0\}$, and so $\mathfrak{I}=\{0\}$ by (1). Finally, $A$ is semiprimitive because $\mathscr{J}(A)\subseteq\mathfrak{I}$. ∎
Proposition 3.10.
Assume that $B=\prod_{i=1}^{t}\mathcal{D}_{i}$, where each $\mathcal{D}_{i}$ is a division ring. For each $1\leq i\leq t$, let $e_{i}=1_{\mathcal{D}_{i}}$ and let $D_{i}=De_{i}\cong D$. Assume also that $A=\prod_{i=1}^{t}A_{i}$, where each $A_{i}$ is a $D_{i}$-subalgebra of $\mathcal{D}_{i}$. If $D$ is one-dimensional and semiprimitive, then the intersection of all the maximal ideals of $A$ is $\{0\}$. Consequently, $A$ is semiprimitive.
Proof.
Let $\mathfrak{I}=\bigcap_{\mathfrak{m}\in\textnormal{Max}(A)}\mathfrak{m}$. Since $A$ is a direct product of rings, we have $\mathfrak{I}=\prod_{i=1}^{t}\mathfrak{I}_{i}$, where each $\mathfrak{I}_{i}$ is a two-sided ideal of $A_{i}$. We claim that
$\mathfrak{I}_{i}=\bigcap_{\mathfrak{m}_{i}\in\textnormal{Max}(A_{i})}\mathfrak{m}_{i}$
for each $i$. To see this, note that each maximal ideal $\mathfrak{m}\subseteq A$ has the form $\mathfrak{m}=\prod_{i=1}^{t}\mathfrak{a}_{i}$, where $\mathfrak{a}_{j}$ is a maximal ideal of $A_{j}$ for exactly one $j\in\{1,\ldots,t\}$, and $\mathfrak{a}_{i}=A_{i}$ for $i\neq j$. Intersecting all of the maximal ideals of $A$ yields the desired decomposition of $\mathfrak{I}$.
Now, if $D$ is semiprimitive, then so is each $D_{i}$. Hence, $\mathfrak{I}_{i}=\{0\}$ for each $1\leq i\leq t$ by Lemma 3.9(2), and therefore $\mathfrak{I}=\{0\}$. As in Proposition 3.9, AA is semiprimitive because $\mathscr{J}(A)\subseteq\mathfrak{I}$. ∎
Theorem 3.11.
If $\textnormal{Int}_{K}(A)$ is Prüfer and $D$ is semiprimitive, then $A\cong\prod_{i=1}^{t}A_{i}$, where $A_{i}$ is the integral closure of $D$ in a finite field extension $F_{i}/K$. In particular, each $A_{i}$ is an ADDB-domain and $A$ is commutative.
Proof.
Assume $\textnormal{Int}_{K}(A)$ is Prüfer. First, we prove that $A$ is a commutative ring. Let $\mathfrak{I}$ be the intersection of all the maximal ideals of $A$. Since $\textnormal{Int}_{K}(A)$ is Prüfer and $D$ is semiprimitive, by Theorem 3.4 and Proposition 3.10, $\mathfrak{I}=\{0\}$. Moreover, by Corollary 3.7, $A/\mathfrak{m}$ is a finite field for each $\mathfrak{m}\in\textnormal{Max}(A)$. So, for all $a,b\in A$ and all $\mathfrak{m}\in\textnormal{Max}(A)$, we have $ab-ba\in\mathfrak{m}$. Hence, $ab-ba\in\mathfrak{I}$, and therefore $ab=ba$.
Next, apply Theorem 3.4 and express $B$ and $A$ as $B\cong\prod_{i=1}^{t}\mathcal{D}_{i}$ and $A\cong\prod_{i=1}^{t}A_{i}$, where each $\mathcal{D}_{i}$ is a division ring and each $A_{i}$ is an integrally closed domain. For each $1\leq i\leq t$, let $e_{i}=1_{\mathcal{D}_{i}}$, $D_{i}=De_{i}\cong D$, and $K_{i}=Ke_{i}\cong K$. Then, $\mathcal{D}_{i}=(D_{i}\setminus\{0\})^{-1}A_{i}$ for each $i$. Since $A$ is commutative, so is each $A_{i}$. Thus, each $\mathcal{D}_{i}$ is a field $F_{i}$, and $F_{i}:K_{i}]$ is finite because $\dim_{K}(B)$ is finite.
Since $\textnormal{Int}(D)$ is Prüfer, by Theorem 1.3, DD—and hence each $D_{i}$—is an ADDB-domain. From Theorem 3.4, we know that $A=A^{\prime}$ and $A_{i}=A_{i}^{\prime}$ for each $i$. This means that $A_{i}$ is the integral closure of $D_{i}$ in $F_{i}$, and is itself an ADDB-domain by [12, Theorem 36.1(a)]. ∎
If $D$ is not semiprimitive, then there exist examples of noncommutative rings $A$ such that $\textnormal{Int}_{K}(A)$ is Prüfer; see Theorem 5.4. However, the following partial converse of Theorem 3.11 is true.
Theorem 3.12.
Let $D$ be an ADDB-domain. Let $A\cong\prod_{i=1}^{t}A_{i}$, where $A_{i}$ is the integral closure of $D$ in a finite field extension $F_{i}/K$. Then $\textnormal{Int}_{K}(A)$ is Prüfer.
Proof.
Note that each $A_{i}$ is an ADDB-domain [12, Theorem 36.1(a)]. Let $F$ be a finite Galois extension of $K$ containing $F_{i}$ for $i=1,\ldots,t$ and let $D_{F}$ be the integral closure of $D$ in $F$. Note that $D_{F}$ is also an ADDB-domain. So, $\textnormal{Int}(D_{F})$ is Prüfer by Theorem 1.3.
Let $G=\text{Gal}(F/K)$, let $i\in\{1,\ldots,t\}$, and let $G(A_{i})=\bigcup_{\sigma\in G}\sigma(A_{i})$. It is easy to show that $G(A_{i})=\Omega_{A_{i}}$ and $\textnormal{Int}_{K}(A_{i})=\textnormal{Int}_{K}(G(A_{i}),\overline{D})=\textnormal{Int}_{K}(G(A_{i}),D_{F})$, and this last ring is integrally closed by Theorem 1.6. Hence, we have
$\textnormal{Int}_{K}(A)=\bigcap_{i=1}^{n}\textnormal{Int}_{K}(A_{i})=\bigcap_{i=1}^{n}\textnormal{Int}_{K}(G(A_{i}),D_{F})=\textnormal{Int}_{K}(\bigcup_{i=1}^{n}G(A_{i}),D_{F}).$
Now, by [26, Lemma 2.20], $\textnormal{Int}_{F}(\bigcup_{i=1}^{n}G(A_{i}),D_{F})$ is the integral closure of $\textnormal{Int}_{K}(\bigcup_{i=1}^{n}G(A_{i}),D_{F})$ in $F(X)$. Since $\textnormal{Int}(D_{F})\subseteq\textnormal{Int}_{F}(\bigcup_{i=1}^{n}G(A_{i}),D_{F})$ and the former ring is a Prüfer domain, it follows that $\textnormal{Int}_{F}(\bigcup_{i=1}^{n}G(A_{i}),D_{F})$ is also Prüfer. By [12, Theorem 22.4], $\textnormal{Int}_{K}(\bigcup_{i=1}^{n}G(A_{i}),D_{F})$ is Prüfer, as we wanted to prove. ∎
Corollary 3.13.
Let $D$ be a semiprimitive ADDB-domain. Then, $\textnormal{Int}_{K}(A)$ is Prüfer if and only if $A\cong\prod_{i=1}^{t}A_{i}$, where $A_{i}$ is the integral closure of $D$ in a finite field extension $F_{i}/K$.
4 Sufficient conditions
In this section, we complete the proof of Theorem 1.7 by showing that if $D$ is an ADDB-domain and $A=A^{\prime}$, then $\textnormal{Int}_{K}(A)$ is Prüfer. This is accomplished by constructing a family of Prüfer domains within $K[X]$, one of which must be a subring of $\textnormal{Int}_{K}(A,A^{\prime}):=\{f\in K[X]\mid f(A)\subseteq A^{\prime}\}$ (which is equal to $\textnormal{Int}_{K}(A)$ if $A=A^{\prime}$). To construct these Prüfer domains, we generalize [19, Section 2], which dealt with a similar problem for $D=\mathbb{Z}$.
Recall from the introduction that $\overline{K}$ denotes a fixed algebraic closure of $K$, and from Definition 1.5 that for each $n\geq 1$, $\Lambda_{n}$ is the set of elements of $\overline{K}$ that are integral over $D$ of degree at most $n$. For $\mathscr{E}\subseteq\Lambda_{n}$, define $\textnormal{Int}_{K}(\mathscr{E},\Lambda_{n})=\{f\in K[X]\mid f(\mathscr{E})\subseteq\Lambda_{n}\}$. As usual, we let $\textnormal{Int}_{K}(\Lambda_{n}):=\textnormal{Int}_{K}(\Lambda_{n},\Lambda_{n})$.
It is shown in [27, Prop. 7] that for any $\mathscr{E}\subseteq\Lambda_{n}$, $\textnormal{Int}_{K}(\mathscr{E},\Lambda_{n})$ is an integrally closed subring of $K[X]$. We wish to prove that $\textnormal{Int}_{K}(\Lambda_{n})$ is a Prüfer domain for each $n\geq 1$. To do this, we provide a broad construction that produces a Prüfer domain within $K[X]$.
4.1 Building a Prüfer domain inside $K[X]$
Construction 4.1.
Let $D$ be an almost Dedekind domain with finite residue fields. As mentioned in Notation 1.1, $D$ contains a principal ideal domain $D_{0}$ and we use $\mathfrak{p}$ to denote a nonzero prime element of $D_{0}$.
For each nonzero prime $P$ of $D$, let $I_{P}$ be an index set and let $\{V_{P,i}\mid i\in I_{P}\}$ be a collection of valuation overrings of $D[X]$ such that for each $i\in I_{P}$, the valuation $v_{P,i}$ of $K(X)$ extends $v_{P}$, and $V_{P,i}$ has maximal ideal $M_{P,i}$. Let $\mathfrak{p}$ be the prime of $D_{0}$ below $P$. Assume that the following two conditions hold:
  1. (i)
    there exists $e_{\mathfrak{p}}\geq 1$ such that for all $P|\mathfrak{p}D_{0}$ and all $i\in I_{P}$, $M_{P,i}^{e_{\mathfrak{p}}}\subseteq\mathfrak{p}V_{P,i}$, and
  2. (ii)
    there exists $f_{\mathfrak{p}}\geq 1$ such that for all $P|\mathfrak{p}D_{0}$ and all $i\in I_{P}$, $(V_{P,i}/M_{P,i}):(D_{0}/\mathfrak{p}D_{0})]\leq f_{\mathfrak{p}}$.
Let $\text{Spec}(D)$ and $\textnormal{Max}(D)$ be, respectively, the sets of prime and maximal ideals of $D$. We then form the following subring of $K[X]$:
$\mathscr{D}:=\bigcap_{P\in\textnormal{Max}(D)}\Big{(}\;\bigcap_{i\in I_{P}}V_{P,i}\Big{)}\cap K[X].$
Remark 4.2.
Since $D$ is assumed to be an almost Dedekind domain, each valuation $v_{P}$ is discrete. The existence of $e_{\mathfrak{p}}$ implies that for each $P|\mathfrak{p}D_{0}$ and each $i\in I_{P}$, there exists a positive integer $\lambda_{P,i}\leq e_{\mathfrak{p}}$ such that $M_{P,i}^{\lambda_{P,i}}=\mathfrak{p}V_{P,i}$. Indeed, the containments $M_{P,i}^{e_{\mathfrak{p}}}\subseteq\mathfrak{p}V_{P,i}\subseteq M_{P,i}$ imply that $\mathfrak{p}V_{P,i}$ is $M_{P,i}$-primary, so by [12, Theorem 17.3(b)] $\mathfrak{p}V_{P,i}$ is equal to a power $M_{P,i}^{\lambda_{P,i}}$ of the maximal ideal $M_{P,i}$ and necessarily $\lambda_{P,i}\leq e_{\mathfrak{p}}$.
For the remainder of this section, assume that $\mathscr{D}$ is a domain formed via Construction 4.1. We will prove that $\mathscr{D}_{\mathfrak{M}}$ is a valuation domain for each maximal ideal $\mathfrak{M}$ of $\mathscr{D}$, and hence that $\mathscr{D}$ is a Prüfer domain. The ideals of $\mathscr{D}$ fall into two types. An ideal $\mathfrak{I}\subseteq\mathscr{D}$ is unitary if $\mathfrak{I}\cap D\neq\{0\}$, and is non-unitary otherwise. The non-unitary maximal ideals of $\mathscr{D}$ can be dealt with quite quickly.
Lemma 4.3.
If $\mathfrak{M}$ is a non-unitary maximal ideal of $\mathscr{D}$, then $\mathscr{D}_{\mathfrak{M}}$ is a valuation domain.
Proof.
When $\mathfrak{M}$ is non-unitary and maximal, we have $K[X]\subseteq\mathscr{D}_{\mathfrak{M}}$. Thus, $\mathscr{D}_{\mathfrak{M}}$ is a local overring of the Prüfer domain $K[X]$, and hence is a valuation domain. ∎
Working with the unitary maximal ideals of $\mathscr{D}$ is more complicated. When $\mathfrak{M}$ is unitary, its contraction to $D$ is a nonzero prime $P\subseteq D$, and $P$ itself lies above a prime element $\mathfrak{p}$ of $D_{0}$.
Notation 4.4.
Recall the definitions of the positive integers $e_{\mathfrak{p}}$ and $f_{\mathfrak{p}}$ from Construction 4.1. For each $\mathfrak{p}$, let $s_{\mathfrak{p}}=e_{\mathfrak{p}}!$ and $r_{\mathfrak{p}}=|D_{0}/\mathfrak{p}D_{0}|^{f_{\mathfrak{p}}!}$. Note that $r_{\mathfrak{p}}\geq 2$.
Proposition 4.5.
  1. (1)
    For each $\mathfrak{p}$ and each $f\in\mathscr{D}$, $\dfrac{(f^{r_{\mathfrak{p}}}-f)^{s_{\mathfrak{p}}}}{\mathfrak{p}}\in\mathscr{D}$.
  2. (2)
    Let $\mathfrak{M}$ be a unitary prime ideal of $\mathscr{D}$ above $\mathfrak{p}$. Then, $f^{r_{\mathfrak{p}}}-f\in\mathfrak{M}$ for all $f\in\mathscr{D}$.
Proof.
For readability, let $r=r_{\mathfrak{p}}$ and $s=s_{\mathfrak{p}}$.
(1) First, we show that $(f^{r}-f)^{s}\in\mathfrak{p}V_{P,i}$ for each $P|\mathfrak{p}D_{0}$ and each $i\in I_{P}$. By construction, $(V_{P,i}/M_{P,i}):(D_{0}/\mathfrak{p}D_{0})]$ divides $f_{\mathfrak{p}}!$ and $e_{\mathfrak{p}}\leq s$. So, $f^{r}-f\in M_{P,i}$, and then $(f^{r}-f)^{s}\in M_{P,i}^{s}\subseteq\mathfrak{p}V_{P,i}$. Thus, $(f^{r}-f)^{s}/\mathfrak{p}\in V_{P,i}$.
Next, if $P^{\prime}$ is a nonzero prime of $D$ that is not above $\mathfrak{p}$, then $\mathfrak{p}$ is invertible in each valuation overring $V^{\prime}$ of $D$ associated to $P^{\prime}$. In this case, $(f^{r}-f)^{s}/\mathfrak{p}\in V^{\prime}$. We conclude that $(f^{r}-f)^{s}/\mathfrak{p}\in\mathscr{D}$.
(2) By part (1), $(f^{r}-f)^{s}\in\mathfrak{p}\mathscr{D}\subseteq\mathfrak{M}$ for each $f\in\mathscr{D}$. Consequently, $f^{r}-f\in\mathfrak{M}$. ∎
Lemma 4.6.
Let $f\in\mathscr{D}$. For every $\mathfrak{p}$, every $P|\mathfrak{p}D_{0}$, and each $i\in I_{P}$, one of the following two conditions holds. Either
  1. (i)
    $v_{P,i}\Big{(}\dfrac{f^{s_{\mathfrak{p}}}}{\mathfrak{p}^{t}}\Big{)}=0$ for some $t\geq 0$, or
  2. (ii)
    $v_{P,i}\Big{(}\dfrac{f^{s_{\mathfrak{p}}}}{\mathfrak{p}^{t}}\Big{)}>0$ for all $t\geq 0$.
Proof.
Fix $\mathfrak{p}$, $P$, and $i$. Let $V=V_{P,i}$, $M=M_{P,i}$, $v=v_{P,i}$, and $s=s_{\mathfrak{p}}$. Since $f\in\mathscr{D}$, $v(f)\geq 0$. If $v(f)=0$, then condition (i) holds with $t=0$. So, assume that $v(f)>0$. This means that $f\in M$. If condition (ii) holds, then we are done. So, assume not. Then, there exists some $n\geq 1$ such that $\mathfrak{p}^{n}V\subseteq f^{s}V$. By the same argument of Remark 4.2, $fV=M^{a}$, for some $a\geq 1$, so $f^{s}V=M^{sa}$. Recall from Remark 4.2 that $\mathfrak{p}V=M^{\lambda}$ for some integer $\lambda\geq 1$. Note that $\lambda$ divides $s$ because $\lambda\leq e_{\mathfrak{p}}$. Hence, there exists some $t\geq 1$ such that $sa=\lambda t$, so that $f^{s}V=M^{\lambda t}=\mathfrak{p}^{t}V$. Hence, $v(f^{s})=v(\mathfrak{p}^{t})$, which is (i). ∎
Lemma 4.7.
Let $f\in\mathscr{D}$. Assume that for some $\mathfrak{p}$, every $P|\mathfrak{p}D_{0}$, and every $i\in I_{P}$, $f$ satisfies one of the following two conditions. Either
  1. (i)
    $v_{P,i}\Big{(}\dfrac{f}{\mathfrak{p}^{t}}\Big{)}=0$ for some $t\geq 0$, or
  2. (ii)
    $v_{P,i}\Big{(}\dfrac{f}{\mathfrak{p}^{t}}\Big{)}>0$ for all $t\geq 0$.
Let $g=\dfrac{f(f^{r_{\mathfrak{p}}-1}-1)^{s_{\mathfrak{p}}}}{\mathfrak{p}}\in K[X]$. Then, $g$ also satisfies (i) or (ii), and $g\in\mathscr{D}$.
Proof.
As before, let $r=r_{\mathfrak{p}}$ and $s=s_{\mathfrak{p}}$. First we show that $g$ satisfies either (i) or (ii) for each $P|\mathfrak{p}D_{0}$ and each $i\in I_{P}$. Fix $P$ and $i\in I_{P}$, let $V=V_{P,i}$, $M=M_{P,i}$, and $v=v_{P,i}$. By Proposition 4.5(2), $f(f^{r-1}-1)=f^{r}-f\in\mathfrak{M}$ for each unitary maximal ideal of $\mathscr{D}$ above $\mathfrak{p}$. So, either $v(f)>0$ or $v(f^{r-1}-1)>0$. Note that these two conditions cannot occur simultaneously.
Assume first that $v(f)>0$. Then, $v(g)=v(\tfrac{f}{\mathfrak{p}})$. If (i) holds for $f$, then $v(f)=tv(\mathfrak{p})$ for some integer $t\geq 1$. So, $v(g)=(t-1)v(\mathfrak{p})$ and (i) holds for $g$. On the other hand, if $f$ satisfies (ii), then so does $g$, because for all $t\geq 0$ we have $v(\tfrac{g}{\mathfrak{p}^{t}})=v(\tfrac{f}{\mathfrak{p}^{t+1}})>0$.
Next, assume that $v(f^{r-1}-1)>0$. Then, $f^{r-1}-1\in M$, $(f^{r-1}-1)^{s}\in\mathfrak{p}V$ and $v(f)=0$ (so, in this case, condition (i) is true for $f$ with $t=0$). By Lemma 4.6, either condition (i) or condition (ii) holds for $(f^{r-1}-1)^{s}$. Moreover, if (i) is true for $(f^{r-1}-1)^{s}$, then the fact that $(f^{r-1}-1)^{s}\in\mathfrak{p}V$ means $v((f^{r-1}-1)^{s}/\mathfrak{p}^{t})=0$ for some $t\geq 1$. We have $v(g)=v((f^{r-1}-1)^{s}/\mathfrak{p})$, so either (i) or (ii) holds for $g$.
Finally, the fact that $g$ satisfies either (i) or (ii) implies that $g\in V$, and this is true for every $P|\mathfrak{p}D_{0}$ and every $i\in I_{P}$. This is enough to conclude that $g\in\mathscr{D}$, because if $P^{\prime}$ is a nonzero prime of $D$ that is not over $\mathfrak{p}$, then $\mathfrak{p}$ is a unit in each valuation overring associated to $P^{\prime}$. ∎
Lemma 4.8.
Let $\mathfrak{M}$ be a unitary prime ideal of $\mathscr{D}$ above $\mathfrak{p}$ and let $f\in\mathscr{D}\setminus\{0\}$. Then, either $f^{s_{\mathfrak{p}}}/\mathfrak{p}^{t}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for all integers $t\geq 1$, or $f^{s_{\mathfrak{p}}}/\mathfrak{p}^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$ for some integer $t\geq 0$.
Proof.
Once again, let $r=r_{\mathfrak{p}}$ and $s=s_{\mathfrak{p}}$. There is nothing to prove if $f\notin\mathfrak{M}$, so assume that $f\in\mathfrak{M}$. Let $f_{0}=f^{s}$, and for each $k\geq 1$, let
$f_{k}=\dfrac{f_{k-1}(f_{k-1}^{r-1}-1)^{s}}{\mathfrak{p}}.$
By Lemma 4.6, $f_{0}$ meets the conditions of Lemma 4.7. Thus, $f_{1}\in\mathscr{D}$, and by induction $f_{k}\in\mathscr{D}$ for all $k\geq 0$.
For each $k\geq 1$, let $h_{k}=\prod_{i=0}^{k-1}(f_{i}^{r-1}-1)^{s}$. Then, $f_{k}=(f^{s}h_{k})/\mathfrak{p}^{k}$ for all $k\geq 1$. Notice that if $f_{i}\in\mathfrak{M}$ for all $0\leq i\leq k-1$, then $h_{k}\notin\mathfrak{M}$. Hence, if there exists a smallest positive integer $t$ such that $f_{t}\notin\mathfrak{M}$, then $f^{s}/\mathfrak{p}^{t}=f_{t}/h_{t}$ is a unit in $\mathscr{D}_{\mathfrak{M}}$. Otherwise, $f_{t}\in\mathfrak{M}$ for every $t\geq 1$, each $h_{t}\notin\mathfrak{M}$, and $f^{s}/\mathfrak{p}^{t}=f_{t}/h_{t}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for all $t\geq 1$. ∎
Theorem 4.9.
Let $\mathscr{D}$ be a domain formed via Construction 4.1. Then, $\mathscr{D}$ is Prüfer.
Proof.
Let $\mathfrak{M}$ be a maximal ideal of $\mathscr{D}$. If $\mathfrak{M}$ is non-unitary, then $\mathscr{D}_{\mathfrak{M}}$ is a valuation domain by Lemma 4.3. So, assume that $\mathfrak{M}$ is a unitary maximal ideal above $\mathfrak{p}$. Let $f,g\in\mathscr{D}$ be nonzero and coprime in $K[X]$. We will show that either $f/g$ or $g/f$ is in $\mathscr{D}_{\mathfrak{M}}$. This is clear if either $f\notin\mathfrak{M}$ or $g\notin\mathfrak{M}$. So, assume that both $f$ and $g$ are in $\mathfrak{M}$.
Let $s=s_{\mathfrak{p}}$. Since $f$ and $g$ are coprime in $K[X]$, there exist $h_{1},h_{2}\in K[X]$ such that $1=h_{1}f^{s}+h_{2}g^{s}$. Choose $d\in D\setminus\{0\}$ such that $dh_{1},dh_{2}\in D[X]\subseteq\mathscr{D}$. If both $f^{s}/\mathfrak{p}^{t}$ and $g^{s}/\mathfrak{p}^{t}$ are in $\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for every positive integer $t$, then
$\dfrac{d}{\mathfrak{p}^{t}}=\dfrac{dh_{1}f^{s}+dh_{2}g^{s}}{\mathfrak{p}^{t}}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$
for all $t\geq 1$. This implies that $d/\mathfrak{p}^{t}\in P=\mathfrak{M}\cap D$ for all $t\geq 1$, which is impossible. By Lemma 4.8, we can find $t\geq 1$ such that both $f^{s}/\mathfrak{p}^{t}$ and $g^{s}/\mathfrak{p}^{t}$ are in $\mathscr{D}_{\mathfrak{M}}$, and either $f^{s}/\mathfrak{p}^{t}$ or $g^{s}/\mathfrak{p}^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$. Without loss of generality, assume that $f^{s}/\mathfrak{p}^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$. Then,
$(g/f)^{s}=g^{s}/f^{s}=(g^{s}/\mathfrak{p}^{t})/(f^{s}/\mathfrak{p}^{t})\in\mathscr{D}_{\mathfrak{M}}.$ (4.10)
Since $\mathscr{D}$ is an intersection of valuation domains, it is integrally closed. So, $\mathscr{D}_{\mathfrak{M}}$ is also integrally closed. Thus, $g/f\in\mathscr{D}_{\mathfrak{M}}$, $\mathscr{D}_{\mathfrak{M}}$ is a valuation domain, and $\mathscr{D}$ is Prüfer. ∎
4.2 Proving that $\textnormal{Int}_{K}(\Lambda_{n})$ is Prufer
Next, we show how to express $\textnormal{Int}_{K}(\Lambda_{n})$ as the intersection of a collection of valuation domains like those of Construction 4.1. It then follows from Theorem 4.9 that $\textnormal{Int}_{K}(\Lambda_{n})$ is Prüfer, and this will be enough to complete the proof of Theorem 1.7(1).
For each $P\in\textnormal{Max}(D)$, we let $\widehat{D_{P}}$ and $\widehat{K}_{P}$ be the $P$-adic completions of $D$ and $K$, respectively. Let $\overline{\widehat{K}_{P}}$ be a fixed algebraic closure of $\widehat{K}_{P}$. Since $D_{P}$ is a rank one valuation domain, it follows that $v_{P}$ extends uniquely to $\overline{\widehat{K}_{P}}$ and the valuation domain of this unique extension is equal to $\overline{\widehat{D_{P}}}$, the absolute integral closure of $\widehat{D_{P}}$. We still denote this unique extension by $v_{P}$. For $n\geq 1$, we set $\widehat{\Lambda}_{P,n}:=\{\alpha\in\overline{\widehat{D_{P}}}\mid[\widehat{K}_{P}(\alpha):\widehat{K}_{P}]\leq n\}$. Finally, for $\alpha\in\overline{\widehat{D_{P}}}$, we define $V_{P,\alpha}:=\{\phi\in K(X)\mid\phi(\alpha)\in\overline{\widehat{D_{P}}}\}$, which is a valuation domain of $K(X)$ lying above $D_{P}$ by [24, Proposition 2.2].
Theorem 4.11.
Let $n\geq 1$. Then,
$\textnormal{Int}_{K}(\Lambda_{n})=\bigcap_{P\in\textnormal{Max}(D)}\bigcap_{\alpha\in\widehat{\Lambda}_{P,n}}V_{P,\alpha}\cap K[X].$
Proof.
Let $f\in\textnormal{Int}_{K}(\Lambda_{n})$ and write $f(X)=\frac{g(X)}{d}$ for some $g\in D[X]$ and $d\in D\setminus\{0\}.$ For each $P\in\textnormal{Max}(D)$ such that $d\notin P$, we have $f\in D_{P}[X]$, so $f\in V_{P,\alpha}$ for each $\alpha\in\widehat{\Lambda}_{P,n}$. Next, let $P\in\textnormal{Max}(D)$ be such that $d\in P$ and fix $\alpha\in\widehat{\Lambda}_{P,n}$. Let $k=v_{P}(d)\in\mathbb{Z}$. Let $p_{\alpha}\in\widehat{D_{P}}[X]$ be the minimal polynomial of $\alpha$.
Extend the valuation $v_{P}$ to $\widehat{K}_{P}[X]$ by taking $v_{P}(h)$ to be the minimum of the values of the coefficients of $h\in\widehat{K}_{P}[X]$. Since $D$ is dense in $\widehat{D_{P}}$, for any $\delta\in\mathbb{Z}$ we can find a monic $q\in D[X]$ such that $\deg(q)=\deg(p_{\alpha})$ and $v_{P}(p_{\alpha}-q)>\delta$. Let $\varepsilon\in\mathbb{Z}$ such that $\varepsilon>\max\{k,v_{P}(\alpha_{i}-\alpha_{j})\mid\alpha_{i}\not=\alpha_{j}\in\Omega_{\alpha}\}$. Then, by [31, Chapter 1, (P)], there exists $q\in D[X]$ monic of degree $\deg(p_{\alpha})$ and a root $\gamma$ of $q$ in $\overline{\widehat{D_{P}}}$ such that $v_{P}(\alpha-\gamma)\geq\varepsilon>k$.
Next, for each root $\lambda\in\Omega_{q}$ of $q$ in $\overline{K}$, let $D_{\lambda}^{\prime}$ be the integral closure of $D$ in $K(\lambda)$. By [31, Chapter 4, (L)], for each $\lambda$ and for each prime $Q$ of $D_{\lambda}^{\prime}$ above $P$, there exists a $K$-embedding $\sigma_{\lambda,Q}:K(\lambda)\to\overline{\widehat{K}_{P}}$ such that the extension $v_{Q}$ of $v_{P}$ to $K(\lambda)$ is given by $v_{Q}=v_{P}\circ\sigma_{\lambda,Q}$. Choose $\beta\in\Omega_{q}$ with $K$-embedding $\sigma:K(\beta)\to\overline{\widehat{K}_{P}}$ such that $\sigma(\beta)$ is equal to the root $\gamma$ of $q$ in $\overline{\widehat{D_{P}}}$ chosen as above. Then, by the previous paragraph, $v_{P}(\alpha-\sigma(\beta))>k$. Since $g$ has coefficients in $D$, $\alpha-\sigma(\beta)$ divides $g(\alpha)-g(\sigma(\beta))$ in $\overline{\widehat{D_{P}}}$. Thus, $v_{P}(g(\alpha)-g(\sigma(\beta)))>k$. Furthermore, $v_{P}(g(\sigma(\beta)))=v_{P}(\sigma(g(\beta)))=v_{Q}(g(\beta))$, and $v_{Q}(g(\beta))\geq k$ because $f\in\textnormal{Int}_{K}(\Lambda_{n})$ and $\beta\in\Lambda_{n}$. It follows that $v_{P}(g(\alpha))\geq k$. Hence, $f(\alpha)=\frac{g(\alpha)}{d}\in\overline{\widehat{D_{P}}}$and $f\in V_{P,\alpha}$.
Conversely, let $f\in K[X]$ be in each of the valuation domains $V_{P,\alpha}$, for $P\in\textnormal{Max}(D)$ and $\alpha\in\widehat{\Lambda}_{P,n}$. Let $\beta\in\Lambda_{n}$ and suppose that $f(X)=\frac{g(X)}{d}$ for some $g\in D[X]$ and $d\in D\setminus\{0\}.$ We set $F=K(\beta)$ and let $D^{\prime}$ be the integral closure of $D$ in $F$. Let $P\in\textnormal{Max}(D)$. If $d\not\in P$, then $f\in D_{P}[X]$. In this case, $f(\beta)\in(D^{\prime})_{Q}$ for each prime ideal $Q$ of $D^{\prime}$ lying above $P$, so $f(\beta)\in(D^{\prime})_{P}$, which is the integral closure of $D_{P}$ in $F$. Suppose now that $d\in P$. Fix a prime ideal $Q$ in $D^{\prime}$ lying above $P$. By [31, Chapter 4, (L)], there is an embedding $\sigma_{Q}:F\to\overline{\widehat{K}_{P}}$ such that the extension $v_{Q}$ of $v_{P}$ to $F$ is given by $v_{Q}=v_{P}\circ\sigma_{Q}$. Let $\alpha=\sigma_{Q}(\beta)$. Since $f\in V_{P,\alpha}$, we have $v_{P}(f(\alpha))\geq 0$. So, $v_{Q}(f(\beta))\geq 0$. This holds for each $Q$ above $P$, so $f(\beta)\in(D^{\prime})_{P}$. We have shown that $f(\beta)$ is in $(D^{\prime})_{P}$ for each $P\in\textnormal{Max}(D)$. It follows that $f(\beta)\in D^{\prime}$ and $f\in\textnormal{Int}_{K}(\Lambda_{n})$. ∎
Corollary 4.12.
Let $D$ be an ADDB-domain.
  1. (1)
    For each $n\geq 1$, $\textnormal{Int}_{K}(\Lambda_{n})$ is Prüfer.
  2. (2)
    Let $A$ be a $D$-algebra that is torsion-free and finitely generated as a $D$-module, and such that $A\cap K=D$. Then, $\textnormal{Int}_{K}(A,A^{\prime})$ is Prüfer.
Proof.
(1) Let $n\geq 1$ and express $\textnormal{Int}_{K}(\Lambda_{n})$ as in Theorem 4.11. Let $\mathfrak{p}$ be a nonzero prime element of $D_{0}$. Since $D$ is an ADDB-domain, we know that the integers $E_{\mathfrak{p}}$ and $F_{\mathfrak{p}}$ from Notation 1.1 are bounded above, say by some integer $d_{\mathfrak{p}}$. Let $P$ be a prime of $D$ above $\mathfrak{p}$. By the results of [24, Proposition 2.2], we have $M_{P,\alpha}^{n}=PV_{P,\alpha}$ and $V_{P,\alpha}/M_{P,\alpha}:D/P]\leq n$ for each $\alpha\in\widehat{\Lambda}_{P,n}$. Therefore, $\textnormal{Int}_{K}(\Lambda_{n})$ is a domain formed by Construction 4.1, because we can take $e_{\mathfrak{p}}=f_{\mathfrak{p}}=nd_{\mathfrak{p}}$. Hence, by Theorem 4.9, $\textnormal{Int}_{K}(\Lambda_{n})$ is a Prüfer domain.
(2) Let $n=\dim_{K}(B)$. Then, $\textnormal{Int}_{K}(A,A^{\prime})=\textnormal{Int}_{K}(\Omega_{A},\Lambda_{n})$ by Theorem 1.6. Since $\Omega_{A}\subseteq\Lambda_{n}$, we have $\textnormal{Int}_{K}(\Lambda_{n})\subseteq\textnormal{Int}_{K}(\Omega_{A},\Lambda_{n})$. By part (1), $\textnormal{Int}_{K}(\Lambda_{n})$ is Prüfer. Hence, $\textnormal{Int}_{K}(A,A^{\prime})$ is also Prüfer. ∎
4.3 The proof of Theorem 1.7
To finish the proof of Theorem 1.7, it remains to consider the conditions that involve localizations at primes of $D$. For a prime $P\subseteq D$, let $(A^{\prime})_{P}=(D\setminus P)^{-1}A^{\prime}$ and $(A_{P})^{\prime}=\{b\in B\mid b\text{ is integral over }D_{P}\}$. It is straightforward to check that $(A^{\prime})_{P}=(A_{P})^{\prime}$, so we may take $A_{P}^{\prime}:=(A^{\prime})_{P}=(A_{P})^{\prime}$ and use the notation $A_{P}^{\prime}$ without ambiguity.
Lemma 4.13.
Let $P\in\text{Spec}(D)$. Then, $\textnormal{Int}_{K}(A)_{P}\subseteq\textnormal{Int}_{K}(A_{P})$ and $\textnormal{Int}_{K}(A,A^{\prime})_{P}\subseteq\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$.
Proof.
We prove that $\textnormal{Int}_{K}(A,A^{\prime})_{P}\subseteq\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$ by employing an induction technique due to Rush [32, Proposition 1.4]. The proof that $\textnormal{Int}_{K}(A)_{P}\subseteq\textnormal{Int}_{K}(A_{P})$ is similar. Let $f\in\textnormal{Int}_{K}(A,A^{\prime})_{P}$ have degree $m$ and let $a/d\in A_{P}$, where $a\in A$ and $d\in D\setminus P$. We use induction on $m$. If $m=0$, then $f\in D_{P}\subseteq A_{P}^{\prime}$ and we are done. Assume that $m\geq 1$ and that any polynomial in $\textnormal{Int}_{K}(A,A^{\prime})_{P}$ of degree at most m1m-1 is an element of $\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$. Let $g(x)=d^{m}f(x)-f(dx)$. Then, both $f(dx)$ and $g$ are in $\textnormal{Int}_{K}(A,A^{\prime})_{P}$, and $\deg g\leq m-1$. By induction, $g\in\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$.
Now, let $R=A_{P}^{\prime}\cap K[a]$. It may the case that $A_{P}^{\prime}$ is not a ring; however, $R$ is equal to the integral closure of $D_{P}$ in $K[a]$, and hence is a ring. Both $g(a/d)$ and $f(a)$ are in $R$, and $d$ is invertible in $R$, so $f(a/d)=d^{-m}(g(a/d)+f(a))\in R\subseteq A_{P}^{\prime}$. This is true for all $a/d\in A_{P}$, so $f\in\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$. ∎
Lemma 4.14.
Let $D$ be an ADDB-domain. The following are equivalent.
  1. (i)
    $\textnormal{Int}_{K}(A)$ is Prüfer.
  2. (ii)
    For each $P\in\text{Spec}(D)$, $\textnormal{Int}_{K}(A)_{P}$ is Prüfer.
  3. (iii)
    For each $P\in\text{Spec}(D)$, $\textnormal{Int}_{K}(A_{P})$ is Prüfer.
  4. (iv)
    For each $P\in\text{Spec}(D)$, $\textnormal{Int}_{K}(A_{P})$ is integrally closed.
  5. (v)
    $\textnormal{Int}_{K}(A)=\textnormal{Int}_{K}(A,A^{\prime})$.
Proof.
(i) $\Rightarrow$ (ii) $\Rightarrow$ (iii) Certainly, $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}_{K}(A)_{P}$ for all $P$, and $\textnormal{Int}_{K}(A)_{P}\subseteq\textnormal{Int}_{K}(A_{P})$ by Lemma 4.13.
(iii) $\Rightarrow$ (iv) This is clear, because Prüfer domains are integrally closed.
(iv) $\Rightarrow$ (v) Assume that $\textnormal{Int}_{K}(A_{P})$ is integrally closed for every $P\in\text{Spec}(D)$. On the one hand, we always have $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}_{K}(A,A^{\prime})$. On the other hand,
$\textnormal{Int}_{K}(A,A^{\prime})\subseteq\bigcap_{P\in\text{Spec}(D)}\textnormal{Int}_{K}(A,A^{\prime})_{P}\subseteq\bigcap_{P\in\text{Spec}(D)}\textnormal{Int}_{K}(A_{P},A_{P}^{\prime}),$
where the second containment is true by Lemma 4.13. By assumption, each $\textnormal{Int}_{K}(A_{P})$ is integrally closed. Since $D$ is an ADDB-domain, $D_{P}$ is a DVR with finite residue field. So, $\textnormal{Int}_{K}(A_{P})=\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$ by Theorem 1.6. This gives
$\textnormal{Int}_{K}(A,A^{\prime})\subseteq\bigcap_{P\in\text{Spec}(D)}\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})=\bigcap_{P\in\text{Spec}(D)}\textnormal{Int}_{K}(A_{P})\subseteq\textnormal{Int}_{K}(A).$
Thus, $\textnormal{Int}_{K}(A)=\textnormal{Int}_{K}(A,A^{\prime})$.
(v) $\Rightarrow$ (i) Apply Corollary 4.12(2). ∎
We now have all the pieces necessary to complete the proof of Theorem 1.7.
Proof of Theorem 1.7.
Part (2) is Corollary 3.13. For part (1), the conditions can be shown to be equivalent as follows.
(i) $\Rightarrow$ (v) This is Theorem 3.4.
(v) $\Rightarrow$ (iii) By assumption, $A_{i}=A_{i}^{\prime}$ for each $i$, so $A=A^{\prime}$.
(iii) $\Rightarrow$ (ii) This is clear.
(ii) $\Rightarrow$ (i) Use Corollary 4.12(2).
(iii) $\Leftrightarrow$ (iv) This is Lemma 3.1.
(i) $\Leftrightarrow$ (vi) $\Leftrightarrow$ (vii) $\Leftrightarrow$ (viii) $\Leftrightarrow$ (ii) This is Lemma 4.14. ∎
We now consider Conjecture 1.8, which states that in order for $\textnormal{Int}_{K}(A)$ to be Prüfer, it is sufficient that $\textnormal{Int}_{K}(A)$ be integrally closed. The conjecture is true at least in the case given in Theorem 1.9.
Proof of Theorem 1.9.
Assume that $\textnormal{Int}_{K}(A)$ is integrally closed and that $\textnormal{Int}_{K}(A)_{P}=\textnormal{Int}_{K}(A_{P})$ for all prime ideals $P$ of $D$. Since $\textnormal{Int}_{K}(A)$ is integrally closed, so is each localization $\textnormal{Int}_{K}(A)_{P}$. But, $\textnormal{Int}_{K}(A_{P})=\textnormal{Int}_{K}(A)_{P}$ and $D_{P}$ is a DVR with finite residue rings, so by Theorem 1.6, $\textnormal{Int}_{K}(A_{P})$ is equal to its integral closure $\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$, which is Prüfer by Corollary 4.12(2). By Lemma 4.14, $\textnormal{Int}_{K}(A)$ is Prüfer. ∎
There exist examples [2, Examples VI.4.15, VI.4.16] of ADDB-domains $D$ that are not Dedekind, and also examples [2, Example VI.4.15] of domains $D$ such that $\textnormal{Int}(D)$ is Prüfer, but does not behave well under localization. Taking $A=D$ shows that neither $D$ being Dedekind nor $\textnormal{Int}_{K}(A)$ being well-behaved under localization is necessary for $\textnormal{Int}_{K}(A)$ to be Prüfer. Consequently, Conjecture 1.8 remains an open problem.
5 A noncommutative example
Here, we give an example of a noncommutative $D$-algebra AA such that $\textnormal{Int}_{K}(A)$ is Prüfer. By Theorem 3.11, $D$ cannot be semiprimitive. Our example uses quaternion algebras over the local domain $\mathbb{Z}_{(2)}$.
Notation 5.1.
Let $D$ be an overring of $\mathbb{Z}$ in $\mathbb{Q}$. The ring of Hurwitz quaternions over $D$ is
[…]
Here, $\mathbf{i}$, $\mathbf{j}$, and $\mathbf{k}$ satisfy $\mathbf{i}^{2}=\mathbf{j}^{2}=-1$ and $\mathbf{i}\mathbf{j}=\mathbf{k}=-\mathbf{j}\mathbf{i}$. The ring $D_{\mathbf{H}}$ also contains the Hurwitz unit $\mathbf{h}=(1+\mathbf{i}+\mathbf{j}+\mathbf{k})/2$, which is a root of $X^{2}-X+1$. The extended algebra $B=D_{\mathbf{H}}\otimes_{D}\mathbb{Q}$ of $D_{\mathbf{H}}$ is the division ring of rational quaternions
$B=\{a_{0}+a_{1}\mathbf{i}+a_{2}\mathbf{j}+a_{3}\mathbf{k}\mid a_{i}\in\mathbb{Q}\text{ for all $i$}\}.$
It is well known that $\alpha=a_{0}+a_{1}\mathbf{i}+a_{2}\mathbf{j}+a_{3}\mathbf{k}\in B$ solves the polynomial $X^{2}-2a_{0}X+N(\alpha)\in\mathbb{Q}[X]$, where $N(\alpha):=a_{0}^{2}+a_{1}^{2}+a_{2}^{2}+a_{3}^{2}$ is the norm of $\alpha$.
Lemma 5.2.
Let $a,b,c,d,n\in\mathbb{Z}$ such that $n\geq 2$. Assume that $a^{2}+b^{2}+c^{2}+d^{2}\equiv 0\mod{4^{n}}$. Then, $a,b,c$, and $d$ are all even.
Proof.
By assumption, $a^{2}+b^{2}+c^{2}+d^{2}\equiv 0\mod{4}$, which implies that $a,b,c$, and $d$ are either all odd or all even. If all four integers are odd, then $a^{2}+b^{2}+c^{2}+d^{2}\equiv 4\mod{8}$, which is a contradiction. So, $a,b,c$, and $d$ are all even. ∎
Lemma 5.3.
Let $D$ be an overring of $\mathbb{Z}$ in $\mathbb{Q}$.
  1. (1)
    If $\alpha\in D_{\mathbf{H}}$, then $N(\alpha)\in D$. Thus, $\alpha$ is integral over $D$.
  2. (2)
    Let $D=\mathbb{Z}_{(2)}$ and let $a_{0},a_{1},a_{2},a_{3},e\in\mathbb{Z}$ be odd. Then, $(a_{0}+a_{1}\mathbf{i}+a_{2}\mathbf{j}+a_{3}\mathbf{k})/(2e)\in D_{\mathbf{H}}$.
Proof.
(1) Let $\alpha\in D_{\mathbf{H}}$. If $\alpha=a_{0}+a_{1}\mathbf{i}+a_{2}\mathbf{j}+a_{3}\mathbf{k}$ with each $a_{i}\in D$, then the result is clear. If not, then each $a_{i}\in D+\tfrac{1}{2}$, which means that $a_{i}=(1+2b_{i})/2$ for some $b_{i}\in D$. Then, for each $i$, $(1+2b_{i})^{2}\in 1+4D$. So,
$(1+2b_{0})^{2}+(1+2b_{1})^{2}+(1+2b_{2})^{2}+(1+2b_{3})^{2}\in 4D,$
and thus
$N(\alpha)=\tfrac{1}{4}\big{(}(1+2b_{0})^{2}+(1+2b_{1})^{2}+(1+2b_{2})^{2}+(1+2b_{3})^{2}\big{)}\in D,$
as required.
(2) For each $0\leq i\leq 3$, let $b_{i}\in\mathbb{Z}$ be such that $a_{i}=2b_{i}+1$. Then, $e$ is a unit of $\mathbb{Z}_{(2)}$, so
$\dfrac{a_{0}+a_{1}\mathbf{i}+a_{2}\mathbf{j}+a_{3}\mathbf{k}}{2e}=e^{-1}(\mathbf{h}+b_{0}+b_{1}\mathbf{i}+b_{2}\mathbf{j}+b_{3}\mathbf{k}),$
which is in $D_{\mathbf{H}}$. ∎
Theorem 5.4.
Let $D=\mathbb{Z}_{(2)}$ and $A=D_{\mathbf{H}}$. Then, $\textnormal{Int}_{\mathbb{Q}}(A)$ is a Prüfer domain.
Proof.
We show that $A=A^{\prime}$. Let $\alpha\in B$ be integral over $D$. We may write
$\alpha=\dfrac{a_{0}+a_{1}\mathbf{i}+a_{2}\mathbf{j}+a_{3}\mathbf{k}}{2^{n}e},$
where $a_{0},a_{1},a_{2},a_{3},n,e\in\mathbb{Z}$, $n\geq 0$, and $e$ is odd. If $n=0$, then $\alpha\in A$. So, assume that $n\geq 1$. Then, we may further assume that at least one $a_{i}$ is odd. Since $\alpha$ is integral over $D$, $N(\alpha)\in D$. Thus, there exist $c,d\in\mathbb{Z}$ with $d$ odd such that
$N(\alpha)=\dfrac{a_{0}^{2}+a_{1}^{2}+a_{2}^{2}+a_{3}^{2}}{2^{2n}e^{2}}=\dfrac{c}{d},$
which is equivalent to
$(a_{0}^{2}+a_{1}^{2}+a_{2}^{2}+a_{3}^{2})d=4^{n}e^{2}c.$
Since $d$ is odd, this means that $a_{0}^{2}+a_{1}^{2}+a_{2}^{2}+a_{3}^{2}\equiv 0\mod{4^{n}}$. If $n\geq 2$, then each $a_{i}$ is even by Lemma 5.2, which is a contradiction. So, $n=1$. Since at least one $a_{i}$ is odd and $a_{0}^{2}+a_{1}^{2}+a_{2}^{2}+a_{3}^{2}\equiv 0\mod{4}$, each $a_{i}$ must be odd. Thus, $\alpha=(a_{0}+a_{1}\mathbf{i}+a_{2}\mathbf{j}+a_{3}\mathbf{k})/(2e)\in A$, because $e$ and all the $a_{i}$ are odd. Therefore, $A=A^{\prime}$, and $\textnormal{Int}_{\mathbb{Q}}(A)$ is Prüfer by Theorem 1.7.

### v2

@@
-A classification of Prüfer domains of integer-valued polynomials on algebras
+accepted for publication in the Bulletin of the London Mathematical Society (2026)
 Abstract
@@
 If $\text{char}(K)=0$, then $D_{0}\cong\mathbb{Z}$.
-If $\text{char}(K)=p$, for some integral prime $p$, then $D_{0}\cong\mathbb{F}_{p}[t]$, the univariate polynomial ring over the finite field $\mathbb{F}_{p}$.
-  • Let $\mathfrak{p}$ denote a nonzero prime element of $D_{0}$.
-Explicitly, if $D_{0}\cong\mathbb{Z}$, then $\mathfrak{p}$ is an integral prime; otherwise, $\mathfrak{p}$ is a polynomial irreducible over $\mathbb{F}_{p}$.
+If $\text{char}(K)$ is positive, then $D_{0}\cong F[t]$, the univariate polynomial ring over the finite field $F$ with $\text{char}(K)$ elements.
+  • Let $p$ denote a nonzero prime element of $D_{0}$. Explicitly, if $D_{0}\cong\mathbb{Z}$, then $p$ is an integral prime; otherwise, $p$ is a polynomial irreducible over $F$.
   • For each prime $P$ of $D$, let $v_{P}$ be the normalized $P$-adic valuation on $K$.
-  • For each $\mathfrak{p}$ and each $P$ over $\mathfrak{p}$, define
-    $\quad e_{P,\mathfrak{p}}=v_{P}(\mathfrak{p})\quad\text{ and }\quad f_{P,\mathfrak{p}}=[(D/P):(D_{0}/\mathfrak{p}D_{0})].$
-  • For each $\mathfrak{p}$, define
-    $E_{\mathfrak{p}}:=\{e_{P,\mathfrak{p}}\mid P|\mathfrak{p}D_{0}\}\quad\text{ and }\quad F_{\mathfrak{p}}:=\{f_{P,\mathfrak{p}}\mid P|\mathfrak{p}D_{0}\}.$
-  • We say that $D$ satisfies the double-boundedness condition if, for each $\mathfrak{p}$, both $E_{\mathfrak{p}}$ and $F_{\mathfrak{p}}$ are bounded.
+  • For each $p$ and each $P$ over $p$, define
+    $\quad e_{P,p}=v_{P}(p)\quad\text{ and }\quad f_{P,p}=[(D/P):(D_{0}/pD_{0})].$
+  • For each $p$, define
+    $E_{p}:=\{e_{P,p}\mid P|pD_{0}\}\quad\text{ and }\quad F_{p}:=\{f_{P,p}\mid P|pD_{0}\}.$
+  • We say that $D$ satisfies the double-boundedness condition if, for each $p$, both $E_{p}$ and $F_{p}$ are bounded.
 Definition 1.2.
@@
 Theorem 2.1.
+Let $D$ be an integrally closed domain.
   1. (1)
@@
 However, we state our intermediary results using only the properties of the integrally closed domain $D$ that are needed to reach the desired conclusion.
+In fact, a number of these results (Lemmas 3.2, 3.5, 3.6, 3.9; Propositions 3.10, 3.11; Theorem 3.7) hold even when $D$ is not integrally closed.
 Recall that an element $b\in B$ is integral over $D$ if $b$ solves a monic polynomial in $D[X]$, and that $A$ is integrally closed if $A=A^{\prime}=\{b\in B\mid b\text{ is integral over }D\}$.
@@
 ∎
-For any ring $R$, let $\mathscr{J}(R)$ be the Jacobson radical of $R$.
+For any ring $R$, let $\mathscr{J}(R)$ be the Jacobson radical of $R$ and let $\textnormal{Max}(R)$ be the set of maximal ideals of $R$.
 We recall that $R$ is semiprimitive if $\mathscr{J}(R)=\{0\}$.
@@
 ∎
-Theorem 3.6.
-Let $D$ be a Prüfer domain and assume that $A$ contains a maximal ideal $\mathfrak{m}$ such that $A/\mathfrak{m}\cong M_{n}(F)$, where $n\geq 2$ and $F$ is a finite field.
-Then, $\textnormal{Int}_{K}(A)$ is not Prüfer.
+Lemma 3.6.
+Let $\mathfrak{m}\in\textnormal{Max}(A)$ and let $P=\mathfrak{m}\cap D$.
+  1. (1)
+    $P$ is a nonzero prime ideal of $D$.
+  2. (2)
+    If $D$ is one-dimensional, then $P\in\textnormal{Max}(D)$.
+  3. (3)
+    If $\textnormal{Int}_{K}(A)$ is Prüfer, then $\in\textnormal{Max}(D), PP$ has finite index in $D$, and $\mathfrak{m}$ has finite index in $A$.
 Proof.
-Let $P=D\cap\mathfrak{m}$.
-We first verify that $P$ is a prime ideal of $D$.
-Certainly, $P$ is a proper ideal of $D$ because $1\notin P$.
+(1) Certainly, $P$ is a proper ideal of $D$ because $1\notin P$.
 Let $cd\in P$.
 Then, $cd\in\mathfrak{m}$, and since both $c$ and $d$ are central in $A$, we have $cAd\subseteq\mathfrak{m}$.
-The ring $A$ is noncommutative, but maximal ideals in noncommutative rings are still prime [16, p.
+The ring $A$ may be noncommutative, but maximal ideals in noncommutative rings are still prime [16, p.
 165], so the condition $cAd\subseteq\mathfrak{m}$ implies that either $c\in\mathfrak{m}$ or $d\in\mathfrak{m}$ [16, Prop.
@@
 Thus, either $c$ or $d$ is in $P$.
-Next, recall that $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}(D)$.
-If $P$ has infinite index in $D$, then by [2, Proposition I.3.4] we have $\textnormal{Int}(D)\subseteq D_{P}[X]$ and so $\textnormal{Int}_{K}(A)$ cannot be Prüfer.
-So, assume $D:P]<\infty$ and consider the $D_{P}$-algebra $A_{P}=(D\setminus P)^{-1}A$.
-Since $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}_{K}(A_{P})$, it suffices to show that $\textnormal{Int}_{K}(A_{P})$ is not Prüfer.
-Because $D$ is Prüfer, $D_{P}$ is a valuation domain.
-If $PD_{P}$ is not principal, then by [2, Proposition I.3.16] we have that $\textnormal{Int}(D_{P})=D_{P}[X]$, so the subring $\textnormal{Int}_{K}(A_{P})$ cannot be Prüfer.
-Suppose now that $PD_{P}$ is principal and let $\pi\in D_{P}$ be a generator of $PD_{P}$ Then, $A_{P}/\pi A_{P}$ has $M_{n}(F)$ as a residue ring and $D_{P}:\pi D_{P}]=[D:P]$ is finite.
-By Lemma 3.5, $\textnormal{Int}_{K}(A_{P})$ is not Prüfer, and we are done.
+Suppose that $P=\{0\}$ and let $R=A/\mathfrak{m}$. Then, $D$ would be isomorphic to a subring of $R$, because
+$D=D/(\mathfrak{m}\cap D)\cong(D+\mathfrak{m})/\mathfrak{m}\subseteq R.$
+Now, $R$ is a simple ring and $D$ is central in $R$, so for each nonzero $d\in D$ we have $dR=R$.
+So, $d^{-1}$ exists in $R$.
+Since $A$ is integral over $D$, this implies that $d^{-1}$ is integral over $D$.
+It follows that $K$ is integral over $D$.
+By [1, Proposition 5.7], $D$ is a field.
+This is a contradiction.
+Thus, $P\neq\{0\}$.
+(2) This follows from (1).
+(3) Assume that $\textnormal{Int}_{K}(A)$ is Prüfer.
+Then, $\textnormal{Int}(D)$ is also Prüfer.
+By [4, Prop.
+6.3], $D$ must be an almost Dedekind domain with finite residue fields.
+By (2), $P\in\textnormal{Max}(D)$, and hence $D/P$ is finite.
+Since $A$ is finitely generated as a $D$-module, $A/\mathfrak{m}$ is also finite.
+∎
+Theorem 3.7.
+Assume that $A$ contains a maximal ideal $\mathfrak{m}$ such that $A/\mathfrak{m}\cong M_{n}(F)$, where $n\geq 2$ and $F$ is a finite field. Then, $\textnormal{Int}_{K}(A)$ is not Prüfer.
+Proof.
+Let $P=D\cap\mathfrak{m}$.
+By Lemma 3.6(1), $P$ is a nonzero prime ideal of $D$.
+Suppose that $\textnormal{Int}_{K}(A)$ is Prüfer.
+Then, as in Lemma 3.6, $D$ is an almost Dedekind domain and $P$ has finite index in $D$.
+So, $PD_{P}$ is principal.
+Let $\pi\in D_{P}$ be a generator of $PD_{P}$.
+Then, $A_{P}/\pi A_{P}$ has $M_{n}(F)$ as a residue ring and $D_{P}:\pi D_{P}]=[D:P]$ is finite.
+By Lemma 3.5(1), there exists $\alpha\in A_{P}$ such that $\alpha^{2}\in\pi^{2}A_{P}$, but $\alpha\notin\pi A_{P}$.
+Let $\alpha=a/d$, where $a\in A$ and $d\in D\setminus P$.
+Then, the ring $\textnormal{Int}_{K}(\{a\},A_{P})$ is not integrally closed, because it contains $X^{2}/(d\pi)^{2}$ but not $X/(d\pi)$.
+So, $\textnormal{Int}_{K}(\{a\},A_{P})$ is not Prüfer, which is a contradiction because $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}_{K}(\{a\},A_{P})$.
 ∎
@@
 Proof.
-Assume that $\textnormal{Int}_{K}(A)$ is Prüfer.
-Then, $\textnormal{Int}_{K}(A)\subseteq\textnormal{Int}(D)$ because $A\cap K=D$, and so $\textnormal{Int}(D)$ is Prüfer.
-By [4, Prop.
-6.3], $D$ must be an almost Dedekind domain with finite residue fields, hence $D$ is a Prüfer domain.
-Since $\textnormal{Int}_{K}(A)$ is Prüfer, the condition of Theorem 3.6 cannot hold.
-Therefore, every simple residue ring of AA must be a finite field.
-∎
+Apply Theorem 3.7 and Lemma 3.6. ∎
 When $D$ is semiprimitive, combining Theorem 3.4 with Corollary 3.8 allows us to show that $A$ is a commutative ring whenever $\textnormal{Int}_{K}(A)$ is Prüfer.
@@
 Towards this end, we now study how $D$ being semiprimitive affects $A$.
-In what follows, for any ring $R$, let $\textnormal{Max}(R)$ be the set of maximal ideals of $R$, called the maximal spectrum of $R$.
 Lemma 3.9.
-Assume that $D$ is one-dimensional.
-  1. (1)
-    Let $\mathfrak{m}\in\textnormal{Max}(A)$. Then, $\mathfrak{m}\cap D\in\textnormal{Max}(D)$.
-  2. (2)
-    Let $M\in\textnormal{Max}(D)$. Then, there exists a maximal ideal of $A$ containing $MA$.
+Assume that $D$ is one-dimensional. Let $M\in\textnormal{Max}(D)$. Then, there exists a maximal ideal of $A$ containing $MA$.
 Proof.
-(1) Let $M=\mathfrak{m}\cap D$.
-Working as in the proof of Theorem 3.6, we see that $M$ is a prime ideal of $D$.
-It remains to show that $M\neq\{0\}$.
-If this were the case, then $D$ would be isomorphic to a subring of $A/\mathfrak{m}$, because
-$D=D/(\mathfrak{m}\cap D)\cong(D+\mathfrak{m})/\mathfrak{m}\subseteq A/\mathfrak{m}.$
-However, $A/\mathfrak{m}$ is a simple ring and $D$ is central in $A$, so this would force $D$ to be a field, which is a contradiction.
-Thus, $M$ is a nonzero prime ideal of the one-dimensional domain $D$, and hence $M$ is maximal.
-(2) Consider the localization $D_{M}$ and the $D_{M}$-algebra $A_{M}$.
+Consider the localization $D_{M}$ and the $D_{M}$-algebra $A_{M}$.
 The ring $A_{M}$ cannot be simple, since then $dA_{M}=A_{M}$ for each $d\in D\setminus\{0\}$.
@@
 So, $A_{M}$ contains a nonzero maximal ideal $\mathfrak{M}$.
-The ideal $\mathfrak{M}\cap D_{M}$ is maximal in $D_{M}$, so we have $MD_{M}=\mathfrak{M}\cap D_{M}$.
+The ideal $\mathfrak{M}\cap D_{M}$ is maximal in $D_{M}$ by Lemma 3.6(2), so we have $MD_{M}=\mathfrak{M}\cap D_{M}$.
 It follows that the contraction of $\mathfrak{M}$ to $A$ is a proper ideal of $A$ that contains $M$.
@@
 Let $D$ be an almost Dedekind domain with finite residue fields.
-As mentioned in Notation 1.1, $D$ contains a principal ideal domain $D_{0}$ and we use $\mathfrak{p}$ to denote a nonzero prime element of $D_{0}$.
+As mentioned in Notation 1.1, $D$ contains a principal ideal domain $D_{0}$ and we use $p$ to denote a nonzero prime element of $D_{0}$.
 For each nonzero prime $P$ of $D$, let $I_{P}$ be an index set and let $\{V_{P,i}\mid i\in I_{P}\}$ be a collection of valuation overrings of $D[X]$ such that for each $i\in I_{P}$, the valuation $v_{P,i}$ of $K(X)$ extends $v_{P}$, and $V_{P,i}$ has maximal ideal $M_{P,i}$.
-Let $\mathfrak{p}$ be the prime of $D_{0}$ below $P$.
+Let $p$ be the prime of $D_{0}$ below $P$.
 Assume that the following two conditions hold:
   1. (i)
-    there exists $e_{\mathfrak{p}}\geq 1$ such that for all $P|\mathfrak{p}D_{0}$ and all $i\in I_{P}$, $M_{P,i}^{e_{\mathfrak{p}}}\subseteq\mathfrak{p}V_{P,i}$, and
+    there exists $e_{p}\geq 1$ such that for all $P|pD_{0}$ and all $i\in I_{P}$, $M_{P,i}^{e_{p}}\subseteq pV_{P,i}$, and
   2. (ii)
-    there exists $f_{\mathfrak{p}}\geq 1$ such that for all $P|\mathfrak{p}D_{0}$ and all $i\in I_{P}$, $(V_{P,i}/M_{P,i}):(D_{0}/\mathfrak{p}D_{0})]\leq f_{\mathfrak{p}}$.
+    there exists $f_{p}\geq 1$ such that for all $P|pD_{0}$ and all $i\in I_{P}$, $(V_{P,i}/M_{P,i}):(D_{0}/pD_{0})]\leq f_{p}$.
 Let $\text{Spec}(D)$ and $\textnormal{Max}(D)$ be, respectively, the sets of prime and maximal ideals of $D$. We then form the following subring of $K[X]$:
@@
 Since $D$ is assumed to be an almost Dedekind domain, each valuation $v_{P}$ is discrete.
-The existence of $e_{\mathfrak{p}}$ implies that for each $P|\mathfrak{p}D_{0}$ and each $i\in I_{P}$, there exists a positive integer $\lambda_{P,i}\leq e_{\mathfrak{p}}$ such that $M_{P,i}^{\lambda_{P,i}}=\mathfrak{p}V_{P,i}$.
-Indeed, the containments $M_{P,i}^{e_{\mathfrak{p}}}\subseteq\mathfrak{p}V_{P,i}\subseteq M_{P,i}$ imply that $\mathfrak{p}V_{P,i}$ is $M_{P,i}$-primary, so by [12, Theorem 17.3(b)] $\mathfrak{p}V_{P,i}$ is equal to a power $M_{P,i}^{\lambda_{P,i}}$ of the maximal ideal $M_{P,i}$ and necessarily $\lambda_{P,i}\leq e_{\mathfrak{p}}$.
+The existence of $e_{p}$ implies that for each $P|pD_{0}$ and each $i\in I_{P}$, there exists a positive integer $\lambda_{P,i}\leq e_{p}$ such that $M_{P,i}^{\lambda_{P,i}}=pV_{P,i}$.
+Indeed, the containments $M_{P,i}^{e_{p}}\subseteq pV_{P,i}\subseteq M_{P,i}$ imply that $pV_{P,i}$ is $M_{P,i}$-primary, so by [12, Theorem 17.3(b)] $pV_{P,i}$ is equal to a power $M_{P,i}^{\lambda_{P,i}}$ of the maximal ideal $M_{P,i}$ and necessarily $\lambda_{P,i}\leq e_{p}$.
 For the remainder of this section, assume that $\mathscr{D}$ is a domain formed via Construction 4.1.
@@
 Working with the unitary maximal ideals of $\mathscr{D}$ is more complicated.
-When $\mathfrak{M}$ is unitary, its contraction to $D$ is a nonzero prime $P\subseteq D$, and $P$ itself lies above a prime element $\mathfrak{p}$ of $D_{0}$.
+When $\mathfrak{M}$ is unitary, its contraction to $D$ is a nonzero prime $P\subseteq D$, and $P$ itself lies above a prime element $p$ of $D_{0}$.
 Notation 4.4.
-Recall the definitions of the positive integers $e_{\mathfrak{p}}$ and $f_{\mathfrak{p}}$ from Construction 4.1.
-For each $\mathfrak{p}$, let $s_{\mathfrak{p}}=e_{\mathfrak{p}}!$ and $r_{\mathfrak{p}}=|D_{0}/\mathfrak{p}D_{0}|^{f_{\mathfrak{p}}!}$.
-Note that $r_{\mathfrak{p}}\geq 2$.
+Recall the definitions of the positive integers $e_{p}$ and $f_{p}$ from Construction 4.1. For each $p$, let $s_{p}=e_{p}!$ and $r_{p}=|D_{0}/pD_{0}|^{f_{p}!}$. Note that $r_{p}\geq 2$.
 Proposition 4.5.
   1. (1)
-    For each $\mathfrak{p}$ and each $f\in\mathscr{D}$, $\dfrac{(f^{r_{\mathfrak{p}}}-f)^{s_{\mathfrak{p}}}}{\mathfrak{p}}\in\mathscr{D}$.
+    For each $p$ and each $f\in\mathscr{D}$, $\dfrac{(f^{r_{p}}-f)^{s_{p}}}{p}\in\mathscr{D}$.
   2. (2)
-    Let $\mathfrak{M}$ be a unitary prime ideal of $\mathscr{D}$ above $\mathfrak{p}$. Then, $f^{r_{\mathfrak{p}}}-f\in\mathfrak{M}$ for all $f\in\mathscr{D}$.
+    Let $\mathfrak{M}$ be a unitary prime ideal of $\mathscr{D}$ above $p$. Then, $f^{r_{p}}-f\in\mathfrak{M}$ for all $f\in\mathscr{D}$.
 Proof.
-For readability, let $r=r_{\mathfrak{p}}$ and $s=s_{\mathfrak{p}}$.
-(1) First, we show that $(f^{r}-f)^{s}\in\mathfrak{p}V_{P,i}$ for each $P|\mathfrak{p}D_{0}$ and each $i\in I_{P}$.
-By construction, $(V_{P,i}/M_{P,i}):(D_{0}/\mathfrak{p}D_{0})]$ divides $f_{\mathfrak{p}}!$ and $e_{\mathfrak{p}}\leq s$.
-So, $f^{r}-f\in M_{P,i}$, and then $(f^{r}-f)^{s}\in M_{P,i}^{s}\subseteq\mathfrak{p}V_{P,i}$.
-Thus, $(f^{r}-f)^{s}/\mathfrak{p}\in V_{P,i}$.
-Next, if $P^{\prime}$ is a nonzero prime of $D$ that is not above $\mathfrak{p}$, then $\mathfrak{p}$ is invertible in each valuation overring $V^{\prime}$ of $D$ associated to $P^{\prime}$.
-In this case, $(f^{r}-f)^{s}/\mathfrak{p}\in V^{\prime}$.
-We conclude that $(f^{r}-f)^{s}/\mathfrak{p}\in\mathscr{D}$.
-(2) By part (1), $(f^{r}-f)^{s}\in\mathfrak{p}\mathscr{D}\subseteq\mathfrak{M}$ for each $f\in\mathscr{D}$. Consequently, $f^{r}-f\in\mathfrak{M}$. ∎
+For readability, let $r=r_{p}$ and $s=s_{p}$.
+(1) First, we show that $(f^{r}-f)^{s}\in pV_{P,i}$ for each $P|pD_{0}$ and each $i\in I_{P}$.
+By construction, $(V_{P,i}/M_{P,i}):(D_{0}/pD_{0})]$ divides $f_{p}!$ and $e_{p}\leq s$.
+So, $f^{r}-f\in M_{P,i}$, and then $(f^{r}-f)^{s}\in M_{P,i}^{s}\subseteq pV_{P,i}$.
+Thus, $(f^{r}-f)^{s}/p\in V_{P,i}$.
+Next, if $P^{\prime}$ is a nonzero prime of $D$ that is not above $p$, then $p$ is invertible in each valuation overring $V^{\prime}$ of $D$ associated to $P^{\prime}$.
+In this case, $(f^{r}-f)^{s}/p\in V^{\prime}$.
+We conclude that $(f^{r}-f)^{s}/p\in\mathscr{D}$.
+(2) By part (1), $(f^{r}-f)^{s}\in p\mathscr{D}\subseteq\mathfrak{M}$ for each $f\in\mathscr{D}$. Consequently, $f^{r}-f\in\mathfrak{M}$. ∎
 Lemma 4.6.
-Let $f\in\mathscr{D}$. For every $\mathfrak{p}$, every $P|\mathfrak{p}D_{0}$, and each $i\in I_{P}$, one of the following two conditions holds. Either
+Let $f\in\mathscr{D}$. For every $p$, every $P|pD_{0}$, and each $i\in I_{P}$, one of the following two conditions holds. Either
   1. (i)
-    $v_{P,i}\Big{(}\dfrac{f^{s_{\mathfrak{p}}}}{\mathfrak{p}^{t}}\Big{)}=0$ for some $t\geq 0$, or
+    $v_{P,i}\Big(\dfrac{f^{s_{p}}}{p^{t}}\Big)=0$ for some $t\geq 0$, or
   2. (ii)
-    $v_{P,i}\Big{(}\dfrac{f^{s_{\mathfrak{p}}}}{\mathfrak{p}^{t}}\Big{)}>0$ for all $t\geq 0$.
+    $v_{P,i}\Big(\dfrac{f^{s_{p}}}{p^{t}}\Big)>0$ for all $t\geq 0$.
 Proof.
-Fix $\mathfrak{p}$, $P$, and $i$.
-Let $V=V_{P,i}$, $M=M_{P,i}$, $v=v_{P,i}$, and $s=s_{\mathfrak{p}}$.
+Fix $p$, $P$, and $i$.
+Let $V=V_{P,i}$, $M=M_{P,i}$, $v=v_{P,i}$, and $s=s_{p}$.
 Since $f\in\mathscr{D}$, $v(f)\geq 0$.
@@
 So, assume not.
-Then, there exists some $n\geq 1$ such that $\mathfrak{p}^{n}V\subseteq f^{s}V$.
+Then, there exists some $n\geq 1$ such that $p^{n}V\subseteq f^{s}V$.
 By the same argument of Remark 4.2, $fV=M^{a}$, for some $a\geq 1$, so $f^{s}V=M^{sa}$.
-Recall from Remark 4.2 that $\mathfrak{p}V=M^{\lambda}$ for some integer $\lambda\geq 1$.
-Note that $\lambda$ divides $s$ because $\lambda\leq e_{\mathfrak{p}}$.
-Hence, there exists some $t\geq 1$ such that $sa=\lambda t$, so that $f^{s}V=M^{\lambda t}=\mathfrak{p}^{t}V$.
-Hence, $v(f^{s})=v(\mathfrak{p}^{t})$, which is (i).
+Recall from Remark 4.2 that $pV=M^{\lambda}$ for some integer $\lambda\geq 1$.
+Note that $\lambda$ divides $s$ because $\lambda\leq e_{p}$.
+Hence, there exists some $t\geq 1$ such that $sa=\lambda t$, so that $f^{s}V=M^{\lambda t}=p^{t}V$.
+Hence, $v(f^{s})=v(p^{t})$, which is (i).
 ∎
 Lemma 4.7.
-Let $f\in\mathscr{D}$. Assume that for some $\mathfrak{p}$, every $P|\mathfrak{p}D_{0}$, and every $i\in I_{P}$, $f$ satisfies one of the following two conditions. Either
+Let $f\in\mathscr{D}$. Assume that for some $p$, every $P|pD_{0}$, and every $i\in I_{P}$, $f$ satisfies one of the following two conditions. Either
   1. (i)
-    $v_{P,i}\Big{(}\dfrac{f}{\mathfrak{p}^{t}}\Big{)}=0$ for some $t\geq 0$, or
+    $v_{P,i}\Big(\dfrac{f}{p^{t}}\Big)=0$ for some $t\geq 0$, or
   2. (ii)
-    $v_{P,i}\Big{(}\dfrac{f}{\mathfrak{p}^{t}}\Big{)}>0$ for all $t\geq 0$.
-Let $g=\dfrac{f(f^{r_{\mathfrak{p}}-1}-1)^{s_{\mathfrak{p}}}}{\mathfrak{p}}\in K[X]$. Then, $g$ also satisfies (i) or (ii), and $g\in\mathscr{D}$.
+    $v_{P,i}\Big(\dfrac{f}{p^{t}}\Big)>0$ for all $t\geq 0$.
+Let $g=\dfrac{f(f^{r_{p}-1}-1)^{s_{p}}}{p}\in K[X]$. Then, $g$ also satisfies (i) or (ii), and $g\in\mathscr{D}$.
 Proof.
-As before, let $r=r_{\mathfrak{p}}$ and $s=s_{\mathfrak{p}}$.
-First we show that $g$ satisfies either (i) or (ii) for each $P|\mathfrak{p}D_{0}$ and each $i\in I_{P}$.
+As before, let $r=r_{p}$ and $s=s_{p}$.
+First we show that $g$ satisfies either (i) or (ii) for each $P|pD_{0}$ and each $i\in I_{P}$.
 Fix $P$ and $i\in I_{P}$, let $V=V_{P,i}$, $M=M_{P,i}$, and $v=v_{P,i}$.
-By Proposition 4.5(2), $f(f^{r-1}-1)=f^{r}-f\in\mathfrak{M}$ for each unitary maximal ideal of $\mathscr{D}$ above $\mathfrak{p}$.
+By Proposition 4.5(2), $f(f^{r-1}-1)=f^{r}-f\in\mathfrak{M}$ for each unitary maximal ideal of $\mathscr{D}$ above $p$.
 So, either $v(f)>0$ or $v(f^{r-1}-1)>0$.
@@
 Assume first that $v(f)>0$.
-Then, $v(g)=v(\tfrac{f}{\mathfrak{p}})$.
-If (i) holds for $f$, then $v(f)=tv(\mathfrak{p})$ for some integer $t\geq 1$.
-So, $v(g)=(t-1)v(\mathfrak{p})$ and (i) holds for $g$.
-On the other hand, if $f$ satisfies (ii), then so does $g$, because for all $t\geq 0$ we have $v(\tfrac{g}{\mathfrak{p}^{t}})=v(\tfrac{f}{\mathfrak{p}^{t+1}})>0$.
+Then, $v(g)=v(\tfrac{f}{p})$.
+If (i) holds for $f$, then $v(f)=tv(p)$ for some integer $t\geq 1$.
+So, v(g)=(t1)v(p)v(g)=(t-1)v(p) and (i) holds for $g$.
+On the other hand, if $f$ satisfies (ii), then so does $g$, because for all $t\geq 0$ we have $v(\tfrac{g}{p^{t}})=v(\tfrac{f}{p^{t+1}})>0$.
 Next, assume that $v(f^{r-1}-1)>0$.
-Then, $f^{r-1}-1\in M$, $(f^{r-1}-1)^{s}\in\mathfrak{p}V$ and $v(f)=0$ (so, in this case, condition (i) is true for $f$ with $t=0$).
+Then, $f^{r-1}-1\in M$, $(f^{r-1}-1)^{s}\in pV$ and $v(f)=0$ (so, in this case, condition (i) is true for $f$ with $t=0$).
 By Lemma 4.6, either condition (i) or condition (ii) holds for $(f^{r-1}-1)^{s}$.
-Moreover, if (i) is true for $(f^{r-1}-1)^{s}$, then the fact that $(f^{r-1}-1)^{s}\in\mathfrak{p}V$ means $v((f^{r-1}-1)^{s}/\mathfrak{p}^{t})=0$ for some $t\geq 1$.
-We have $v(g)=v((f^{r-1}-1)^{s}/\mathfrak{p})$, so either (i) or (ii) holds for $g$.
-Finally, the fact that $g$ satisfies either (i) or (ii) implies that $g\in V$, and this is true for every $P|\mathfrak{p}D_{0}$ and every $i\in I_{P}$.
-This is enough to conclude that $g\in\mathscr{D}$, because if $P^{\prime}$ is a nonzero prime of $D$ that is not over $\mathfrak{p}$, then $\mathfrak{p}$ is a unit in each valuation overring associated to $P^{\prime}$.
+Moreover, if (i) is true for $(f^{r-1}-1)^{s}$, then the fact that $(f^{r-1}-1)^{s}\in pV$ means $v((f^{r-1}-1)^{s}/p^{t})=0$ for some $t\geq 1$.
+We have $v(g)=v((f^{r-1}-1)^{s}/p)$, so either (i) or (ii) holds for $g$.
+Finally, the fact that $g$ satisfies either (i) or (ii) implies that $g\in V$, and this is true for every $P|pD_{0}$ and every $i\in I_{P}$.
+This is enough to conclude that $g\in\mathscr{D}$, because if $P^{\prime}$ is a nonzero prime of $D$ that is not over $p$, then $p$ is a unit in each valuation overring associated to $P^{\prime}$.
 ∎
 Lemma 4.8.
-Let $\mathfrak{M}$ be a unitary prime ideal of $\mathscr{D}$ above $\mathfrak{p}$ and let $f\in\mathscr{D}\setminus\{0\}$.
-Then, either $f^{s_{\mathfrak{p}}}/\mathfrak{p}^{t}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for all integers $t\geq 1$, or $f^{s_{\mathfrak{p}}}/\mathfrak{p}^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$ for some integer $t\geq 0$.
+Let $\mathfrak{M}$ be a unitary prime ideal of $\mathscr{D}$ above $p$ and let $f\in\mathscr{D}\setminus\{0\}$.
+Then, either $f^{s_{p}}/p^{t}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for all integers $t\geq 1$, or $f^{s_{p}}/p^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$ for some integer $t\geq 0$.
 Proof.
-Once again, let $r=r_{\mathfrak{p}}$ and $s=s_{\mathfrak{p}}$. There is nothing to prove if $f\notin\mathfrak{M}$, so assume that $f\in\mathfrak{M}$. Let $f_{0}=f^{s}$, and for each $k\geq 1$, let
-$f_{k}=\dfrac{f_{k-1}(f_{k-1}^{r-1}-1)^{s}}{\mathfrak{p}}.$
+Once again, let $r=r_{p}$ and $s=s_{p}$. There is nothing to prove if $f\notin\mathfrak{M}$, so assume that $f\in\mathfrak{M}$. Let $f_{0}=f^{s}$, and for each $k\geq 1$, let
+$f_{k}=\dfrac{f_{k-1}(f_{k-1}^{r-1}-1)^{s}}{p}.$
 By Lemma 4.6, $f_{0}$ meets the conditions of Lemma 4.7. Thus, $f_{1}\in\mathscr{D}$, and by induction $f_{k}\in\mathscr{D}$ for all $k\geq 0$.
 For each $k\geq 1$, let $h_{k}=\prod_{i=0}^{k-1}(f_{i}^{r-1}-1)^{s}$.
-Then, $f_{k}=(f^{s}h_{k})/\mathfrak{p}^{k}$ for all $k\geq 1$.
+Then, $f_{k}=(f^{s}h_{k})/p^{k}$ for all $k\geq 1$.
 Notice that if $f_{i}\in\mathfrak{M}$ for all $0\leq i\leq k-1$, then $h_{k}\notin\mathfrak{M}$.
-Hence, if there exists a smallest positive integer $t$ such that $f_{t}\notin\mathfrak{M}$, then $f^{s}/\mathfrak{p}^{t}=f_{t}/h_{t}$ is a unit in $\mathscr{D}_{\mathfrak{M}}$.
-Otherwise, $f_{t}\in\mathfrak{M}$ for every $t\geq 1$, each $h_{t}\notin\mathfrak{M}$, and $f^{s}/\mathfrak{p}^{t}=f_{t}/h_{t}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for all $t\geq 1$.
+Hence, if there exists a smallest positive integer $t$ such that $f_{t}\notin\mathfrak{M}$, then $f^{s}/p^{t}=f_{t}/h_{t}$ is a unit in $\mathscr{D}_{\mathfrak{M}}$.
+Otherwise, $f_{t}\in\mathfrak{M}$ for every $t\geq 1$, each $h_{t}\notin\mathfrak{M}$, and $f^{s}/p^{t}=f_{t}/h_{t}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for all $t\geq 1$.
 ∎
@@
 If $\mathfrak{M}$ is non-unitary, then $\mathscr{D}_{\mathfrak{M}}$ is a valuation domain by Lemma 4.3.
-So, assume that $\mathfrak{M}$ is a unitary maximal ideal above $\mathfrak{p}$.
+So, assume that $\mathfrak{M}$ is a unitary maximal ideal above $p$.
 Let $f,g\in\mathscr{D}$ be nonzero and coprime in $K[X]$.
@@
 So, assume that both $f$ and $g$ are in $\mathfrak{M}$.
-Let $s=s_{\mathfrak{p}}$.
+Let $s=s_{p}$.
 Since $f$ and $g$ are coprime in $K[X]$, there exist $h_{1},h_{2}\in K[X]$ such that $1=h_{1}f^{s}+h_{2}g^{s}$.
 Choose $d\in D\setminus\{0\}$ such that $dh_{1},dh_{2}\in D[X]\subseteq\mathscr{D}$.
-If both $f^{s}/\mathfrak{p}^{t}$ and $g^{s}/\mathfrak{p}^{t}$ are in $\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for every positive integer $t$, then
-$\dfrac{d}{\mathfrak{p}^{t}}=\dfrac{dh_{1}f^{s}+dh_{2}g^{s}}{\mathfrak{p}^{t}}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$
+If both $f^{s}/p^{t}$ and $g^{s}/p^{t}$ are in $\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$ for every positive integer $t$, then
+$\dfrac{d}{p^{t}}=\dfrac{dh_{1}f^{s}+dh_{2}g^{s}}{p^{t}}\in\mathfrak{M}\mathscr{D}_{\mathfrak{M}}$
 for all $t\geq 1$.
-This implies that $d/\mathfrak{p}^{t}\in P=\mathfrak{M}\cap D$ for all $t\geq 1$, which is impossible.
-By Lemma 4.8, we can find $t\geq 1$ such that both $f^{s}/\mathfrak{p}^{t}$ and $g^{s}/\mathfrak{p}^{t}$ are in $\mathscr{D}_{\mathfrak{M}}$, and either $f^{s}/\mathfrak{p}^{t}$ or $g^{s}/\mathfrak{p}^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$.
-Without loss of generality, assume that $f^{s}/\mathfrak{p}^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$.
+This implies that $d/p^{t}\in P=\mathfrak{M}\cap D$ for all $t\geq 1$, which is impossible.
+By Lemma 4.8, we can find $t\geq 1$ such that both $f^{s}/p^{t}$ and $g^{s}/p^{t}$ are in $\mathscr{D}_{\mathfrak{M}}$, and either $f^{s}/p^{t}$ or $g^{s}/p^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$.
+Without loss of generality, assume that $f^{s}/p^{t}$ is a unit of $\mathscr{D}_{\mathfrak{M}}$.
 Then,
-$(g/f)^{s}=g^{s}/f^{s}=(g^{s}/\mathfrak{p}^{t})/(f^{s}/\mathfrak{p}^{t})\in\mathscr{D}_{\mathfrak{M}}.$ (4.10)
+$(g/f)^{s}=g^{s}/f^{s}=(g^{s}/p^{t})/(f^{s}/p^{t})\in\mathscr{D}_{\mathfrak{M}}.$ (4.10)
 Since $\mathscr{D}$ is an intersection of valuation domains, it is integrally closed.
@@
 (1) Let $n\geq 1$ and express $\textnormal{Int}_{K}(\Lambda_{n})$ as in Theorem 4.11.
-Let $\mathfrak{p}$ be a nonzero prime element of $D_{0}$.
-Since $D$ is an ADDB-domain, we know that the integers $E_{\mathfrak{p}}$ and $F_{\mathfrak{p}}$ from Notation 1.1 are bounded above, say by some integer $d_{\mathfrak{p}}$.
-Let $P$ be a prime of $D$ above $\mathfrak{p}$.
-By the results of [24, Proposition 2.2], we have $M_{P,\alpha}^{n}=PV_{P,\alpha}$ and $V_{P,\alpha}/M_{P,\alpha}:D/P]\leq n$ for each $\alpha\in\widehat{\Lambda}_{P,n}$.
-Therefore, $\textnormal{Int}_{K}(\Lambda_{n})$ is a domain formed by Construction 4.1, because we can take $e_{\mathfrak{p}}=f_{\mathfrak{p}}=nd_{\mathfrak{p}}$.
+Let $p$ be a nonzero prime element of $D_{0}$.
+Since $D$ is an ADDB-domain, we know that the integers $E_{p}$ and $F_{p}$ from Notation 1.1 are bounded above, say by some integer $d_{p}$.
+Let $P$ be a prime of $D$ above $p$.
+By the results of [24, Proposition 2.2], we have $M_{P,\alpha}^{e_{P,\alpha}}=PV_{P,\alpha}$ for some $e_{P,\alpha}\leq n$ and $V_{P,\alpha}/M_{P,\alpha}:D/P]\leq n$ for each $\alpha\in\widehat{\Lambda}_{P,n}$.
+Therefore, $\textnormal{Int}_{K}(\Lambda_{n})$ is a domain formed by Construction 4.1, because we can take $e_{p}=f_{p}=nd_{p}$.
 Hence, by Theorem 4.9, $\textnormal{Int}_{K}(\Lambda_{n})$ is a Prüfer domain.
@@
 The proof that $\textnormal{Int}_{K}(A)_{P}\subseteq\textnormal{Int}_{K}(A_{P})$ is similar.
-Let $f\in\textnormal{Int}_{K}(A,A^{\prime})_{P}$ have degree $m$ and let $a/d\in A_{P}$, where $a\in A$ and $d\in D\setminus P$.
+Let $f\in\textnormal{Int}_{K}(A,A^{\prime})_{P}$ have degree $m$.
 We use induction on $m$.
-If $m=0$, then $f\in D_{P}\subseteq A_{P}^{\prime}$ and we are done.
+If $m=0$, then $f\in A_{P}^{\prime}$ and we are done.
 Assume that $m\geq 1$ and that any polynomial in $\textnormal{Int}_{K}(A,A^{\prime})_{P}$ of degree at most m1m-1 is an element of $\textnormal{Int}_{K}(A_{P},A_{P}^{\prime})$.
+Let $a/d\in A_{P}$, where $a\in A$ and $d\in D\setminus P$.
 Let $g(x)=d^{m}f(x)-f(dx)$.